### v1

# HippoMM: Hippocampal-inspired Multimodal Memory for Long Audiovisual Event Understanding

Abstract
Comprehending extended audiovisual experiences remains a fundamental challenge for computational systems. Current approaches struggle with temporal integration and cross-modal associations that humans accomplish effortlessly through hippocampal-cortical networks. We introduce HippoMM, a biologically-inspired architecture that transforms hippocampal mechanisms into computational advantages for multimodal understanding. HippoMM implements three key innovations: (i) hippocampus-inspired pattern separation and completion specifically designed for continuous audiovisual streams, (ii) short-to-long term memory consolidation that transforms perceptual details into semantic abstractions, and (iii) cross-modal associative retrieval pathways enabling modality-crossing queries. Unlike existing retrieval systems with static indexing schemes, HippoMM dynamically forms integrated episodic representations through adaptive temporal segmentation and dual-process memory encoding. Evaluations on our challenging HippoVlog benchmark demonstrate that HippoMM significantly outperforms state-of-the-art approaches (78.2% vs. 64.2% accuracy) while providing substantially faster response times (20.4s vs. 112.5s). Our results demonstrate that translating neuroscientific memory principles into computational architectures provides a promising foundation for next-generation multimodal understanding systems. The code and benchmark dataset are publicly available at https://github.com/linyueqian/HippoMM.
1 Introduction
Imagine hearing a familiar melody through an open window, instantly recalling not merely the music but the complete concert experience: the interplay of stage lights, the energy of the crowd, and even the aroma of summer rain. This remarkable cognitive capability, known as multimodal pattern completion, involves reconstructing integrated experiences from partial, cross-modal stimuli and remains a significant frontier in computational modeling. Although current methods proficiently process static multimodal data [radford2021learning, girdhar2023imagebind] or retrieve relevant segments from extensive video corpora [ren2024videorag], they struggle fundamentally with long-form multimodal comprehension. Specifically, these approaches exhibit limitations in three key dimensions: segmenting continuous audiovisual streams into coherent episodic units, reconstructing rich experiences by integrating cross-modal and temporal information (particularly from partial cues), and flexibly retrieving information across modalities—such as recalling visual details from auditory cues.
In contrast, the human hippocampus robustly achieves these capabilities. As illustrated conceptually in Figure 1 (top), the hippocampus integrates multisensory inputs (e.g., visual and auditory streams) through structures including the entorhinal cortex, organizes continuous experiences into discrete episodes [baldassano2017discovering], employs pattern separation mechanisms typically attributed to the Dentate Gyrus (DG), and performs pattern completion associated primarily with CA3 circuits to store and recall distinct memory traces [yassa2011pattern, rolls2013quantitative]. These processes ultimately yield cohesive episodic representations, mediated by hippocampal circuitry including the CA1 region [eichenbaum2014time].
Figure 1: Conceptual overview of hippocampal versus HippoMM multimodal processing. (Top) Biological hippocampus integrates visual and auditory modalities through the entorhinal cortex and associated circuits (DG, CA3, CA1) to form and recall episodic memories. (Bottom) Proposed HippoMM architecture processes multimodal inputs (video, audio, cross-modal) inspired by hippocampal principles for episodic memory formation and cue-driven retrieval.
Despite this clear biological blueprint, existing computational memory systems often mimic only isolated aspects. They typically operate unimodally [gutierrez2024hipporag, graves2016hybrid] or function as passive retrieval systems [lewis2020retrieval, ren2024videorag], falling short of replicating the active, reconstructive, and crossmodal associative nature of hippocampal memory, particularly for extended audiovisual data streams. To address these limitations, we introduce HippoMM, a novel architecture designed to computationally implement functional principles inspired by hippocampal mechanisms for long form multimodal content understanding, as conceptualized in Figure 1 (bottom). Processing video, audio, and jointly learned multimodal representations, HippoMM aims to emulate the brain’s ability to form and query rich, integrated multimodal memories. While prior work has drawn inspiration from the hippocampus for unimodal tasks or specific memory functions, HippoMM represents, to our knowledge, the first architecture to systematically adapt and integrate a suite of hippocampal memory mechanisms to address the unique challenges of long-form audiovisual understanding. Our contributions are as follows:
  1. 1.
    Hippocampal Dynamics for Long AV Streams: We operationalize dentate gyrus-inspired pattern separation and CA3-like pattern completion specifically for continuous audio-visual data. This allows HippoMM to segment lengthy streams into distinct episodes and reconstruct coherent multimodal experiences even from fragmented cues, moving beyond passive retrieval.
  2. 2.
    Short-to-Long Term Memory Consolidation: We model the transition from detailed perceptual traces (short-term) to abstract semantic representations (long-term) through a neuro-inspired consolidation and semantic replay process. This contrasts with models that use fixed chunking or lack mechanisms for evolving representations over time, enabling more efficient and meaningful long-term storage.
  3. 3.
    Cross-Modal Associative Querying: HippoMM introduces retrieval pathways that explicitly support querying information in one modality using cues from another (e.g., finding what was seen when a specific sound occurred, or vice-versa). This capability, absent in prior multimodal systems that often disregard audio richness or lack such cross-modal retrieval tests, mimics a key aspect of human associative recall. Unlike existing approaches that treat modalities in isolation, our system maintains the temporal co-occurrence relationships essential for reconstructing coherent multimodal experiences from partial cues.
Evaluated on our proposed HippoVlog benchmark, comprising long-form vlogs with challenging cross-modal and unimodal question-answering tasks, HippoMM achieves state-of-the-art accuracy (78.2% average accuracy, significantly outperforming VideoRAG [ren2024videorag]), particularly excelling in cross-modal and visual reasoning tasks. Ablation studies underscore the critical roles of biologically inspired components: removing detailed recall mechanisms reduces visual accuracy by 6%, and eliminating adaptive reasoning results in a 1.6% decline in semantic understanding. The remainder of the paper is structured as follows: Section 2 discusses related research on hippocampal mechanisms and multimodal understanding; Section 3 formalizes the multimodal memory-augmented retrieval task and introduces the HippoVlog benchmark; Section 4 describes the HippoMM architecture in detail; Section 5 presents comprehensive experimental results; and finally, we conclude with discussions on implications and future directions in Section 6. Our work demonstrates that integrating hippocampal principles can advance computational understanding of complex, continuous multimodal experiences, offering a robust foundation toward developing systems with human-like memory and reasoning capabilities.
2 Related Work
2.1 Hippocampal Memory Mechanisms
Neuroscientific research highlights the hippocampus’s crucial ability to form, organize, and retrieve episodic memories. Foundational work established key principles including pattern separation through sparse coding in the dentate gyrus [yassa2011pattern], autoassociative pattern completion in CA3 [rolls2013quantitative], and theta-gamma oscillations for temporal coding [lisman2009prediction]. Recent findings reveal how the hippocampus segments experiences at event boundaries [baldassano2017discovering, ben2013hippocampal], creating hierarchical memory structures [zacks2007event] that optimize storage and recall. Computational models have translated these principles to machine learning through temporal chunking [dubrow2013influence] and neural memory networks [kumaran2016learning], but primarily in unimodal contexts. The hippocampus’s role as a multimodal integrator [teyler2007hippocampal, eichenbaum2014time] provides critical insights for artificial memory systems. Neurobiological studies demonstrate its capacity to bind cross-modal features while maintaining modality-specific details [quiroga2009explicit, mormann2008latency], a capability existing computational models lack. Critically, effectively integrating continuous, cross-modal audiovisual streams remains underdeveloped, even in recent memory architectures [whittington2020tolman], presenting an important research direction.
2.2 Audio-Visual Multimodal Understanding
Modern multimodal systems leverage joint embedding spaces [radford2021learning, girdhar2023imagebind] to align audio-visual modalities, while video-language models [alayrac2022flamingo, zhang2023video] employ temporal attention for sequential processing. Though effective for static retrieval, these methods lack mechanisms for temporal memory organization critical for long-form event understanding. Recent audio-visual models [shu2023audio, guo2025aligned] enhance modality integration through specialized tokens or multi-scale adapters, yet struggle with extended temporal contexts and cross-modal associative memory. Retrieval-augmented approaches [ren2024videorag, yuan2025memory] address long-context challenges through graph-based grounding and cognitive memory steps, but remain biologically uninspired. Cross-modal retrieval frameworks [lin2024mm] mitigate modality bias yet lack temporal organization, while efficiency-focused methods [lin2025speechprune, liu2025keyframe] prune redundant elements without enabling progressive retrieval pathways.
2.3 Biomimetic Memory Architectures
Memory-augmented neural networks have evolved from early differentiable memory systems [santoro2016meta] to sophisticated architectures explicitly mirroring hippocampal-cortical interactions [whittington2020tolman]. Hippocampus-inspired models like HippoRAG [gutierrez2024hipporag] adapt principles of hippocampal indexing theory for knowledge integration, drawing on biological mechanisms similar to the dentate gyrus’ pattern separation [yassa2011pattern] and CA3’s autoassociative recall [rolls2013quantitative] for efficient document retrieval. Similarly, Differentiable Neural Computers [graves2016hybrid] emulate the hippocampus’ sparse temporal coding mechanisms [lisman2009prediction]. However, these approaches remain predominantly unimodal, neglecting the hippocampus’ critical role as a multimodal binding engine that coordinates cortical sensory streams through theta-phase coding, grid cell-inspired spatial organization, and cross-modal integration [eichenbaum2014time].
3 Multimodal Memory-augmented Retrieval
3.1 Task Formulation
We formalize the Multimodal Memory-augmented Retrieval (MMR) task as the challenge of developing systems that process, understand, and reason about long-form continuous multimodal data streams $X=\{x_{t}\}$, exemplified by extended audiovisual recordings. Given a user query $q$ targeting visual ($q_{V}$), auditory ($q_{A}$), cross-modal ($q_{V\times A}$), or semantic ($q_{S}$) information, the system must construct and utilize appropriate memory representations to synthesize a coherent and accurate answer $a$. The task decomposes into three fundamental stages. First, a memory formation function $\mathcal{F}_{\text{mem}}$ transforms the raw input stream $X$ into a structured memory representation $M$:
$M=\mathcal{F}_{\text{mem}}(X)$ (1)
This memory $M=\{m_{k}\}$ comprises encoded segments $m_{k}$, each capturing multimodal features, temporal context, and semantic abstractions. Second, a retrieval function $\mathcal{R}_{\text{retrieval}}$ processes the query $q$ against memory $M$ to identify relevant evidence $E$:
$E=\mathcal{R}_{\text{retrieval}}(q,M)$ (2)
Here, $E$ typically includes a relevance-ranked subset of memory elements (i.e., $E=\{\langle m_{k},r_{k}\rangle\mid m_{k}\in M\land\text{relevant}(m_{k},q)\}$). Finally, an answer synthesis function $\mathcal{G}_{\text{synth}}$ generates the response $a$ by reasoning over the query $q$ and retrieved evidence $E$:
$a=\mathcal{G}_{\text{synth}}(q,E)$ (3)
Implementing an effective MMR system requires sophisticated realizations of $\mathcal{F}_{\text{mem}}$, $\mathcal{R}_{\text{retrieval}}$, and $\mathcal{G}_{\text{synth}}$ that handle temporal dependencies, enable cross-modal associations, and support robust reasoning, analogous to biological memory processes described in cognitive neuroscience [yassa2011pattern, eichenbaum2014time].
3.2 HippoVlog Dataset
To evaluate MMR systems under realistic conditions, we introduce HippoVlog. Existing benchmarks for multimodal understanding typically utilize short video clips, rely solely on Large Language Model (LLM)-based evaluation without ground truth answers [ren2024videorag], or lack audio elements for cross-modal reasoning within continuous, long-form videos [MLVU]. HippoVlog addresses these gaps, comprising 25 long-form daily vlogs (682 minutes total) with naturalistic audiovisual content. It includes 1,000 validated multiple-choice question-answer pairs, each with four candidate answers, targeting specific memory functions:
  • Cross-modal binding ($T_{V\times A}$): Linking visual/auditory cues (e.g., "What object appeared when the speaker mentioned hiking?").
  • Auditory-focused retrieval ($T_{A}$): Extracting audio details (e.g., "What instrument played during the introduction?").
  • Visual-focused retrieval ($T_{V}$): Extracting visual details (e.g., "What text was visible on the store sign?").
  • Semantic reasoning ($T_{S}$): Integrating semantic information over time (e.g., "What are the cooking steps mentioned throughout the video?").
Questions were created using GPT-4o [hurst2024gpt], manually validated, and annotated with timestamps. HippoVlog’s objective assessment format, focus on audio-visual integration, and long-form temporal dynamics make it particularly suitable for evaluating MMR systems. Construction details are provided in the Appendix and our dataset is publicly available [^1] https://github.com/linyueqian/HippoVlog/.
Figure 2: The HippoMM architecture for multimodal memory. (a) Memory Formation: composed of Temporal Pattern Separation ($\mathcal{S}_{t}$) based on perceptual boundaries, Perceptual Encoding of visual and auditory inputs with cross-modal features, Memory Consolidation using similarity-based filtering ($K$), and Semantic Replay generating ThetaEvent representations ($\theta$). (b) Memory Retrieval: implementing Query-Driven Pattern Completion through fast retrieval ($\Phi_{\text{fast}}$) and detailed recall pathways with temporal window localization ($\mathbf{W}$).
4 HippoMM System
4.1 Overview
We introduce HippoMM, a computational architecture designed for the complex task of understanding and reasoning about long-form multimodal data streams, typified by extended audiovisual recordings. The architecture draws functional inspiration from the mammalian hippocampus’s established roles in episodic memory processing [eichenbaum2017prefrontal], including cross-modal binding, temporal sequence organization, pattern separation (differentiating similar inputs), pattern completion (retrieving full memories from partial cues), and memory consolidation. While HippoMM does not model neural dynamics directly, it seeks to replicate key functional principles through distinct algorithmic modules. Our objective is to overcome the limitations of conventional multimodal systems, which often lack robust mechanisms for handling continuous temporal dynamics and associative cross-modal recall. HippoMM enables context-sensitive retrieval, facilitating the reconstruction of integrated multimodal experiences from potentially incomplete queries (e.g., retrieving visual context associated with a specific sound event). The system architecture, depicted schematically in Figure 2, comprises two primary phases: Memory Formation (Sec. 4.2), which processes raw inputs into structured, abstracted representations in long-term memory (Figure 2a), and Memory Retrieval (Sec. 4.3), which leverages these representations to answer user queries via hierarchical and associative search strategies (Figure 2b).
4.2 Memory Formation
4.2.1 Temporal Pattern Separation
Continuous experience is often perceived and remembered as discrete events [baldassano2017discovering]. To computationally mimic the hippocampus’s ability to segment experience (a function often linked to the dentate gyrus [yassa2011pattern]), HippoMM first employs content-adaptive temporal segmentation as a functional approximation. This stage analyzes the incoming raw audiovisual stream to identify boundaries corresponding to significant perceptual shifts, partitioning the stream into potential event segments. This contrasts with fixed-duration chunking, aiming instead for more semantically coherent and efficiently processed units. A segment boundary is indicated at time $t$ by $\mathcal{S}_{t}=1$ if a significant change is detected in either the visual or auditory stream, using computationally efficient metrics applied directly to the input:
$\mathcal{S}_{t}=\bigvee_{m\in\{v,a\}}\text{1}\left[d_{m}(\text{Input}_{m}(t),\text{Input}_{m}(t-1))>\tau_{m}\right]$ (4)
where $d_{m}$ is a modality-specific distance function comparing relevant inputs for modality $m$ at times $t$ and $t-1$, $\tau_{m}$ is a predefined threshold, $\text{1}[\cdot]$ is the indicator function, and $\bigvee$ denotes logical OR. Specifically, for the visual modality ($m=v$), we define the distance $d_{v}$ based on the Structural Similarity Index Measure (SSIM) between consecutive raw visual frames $F_{t},F_{t-1}$:
$d_{v}(F_{t},F_{t-1})=1-\text{SSIM}(F_{t},F_{t-1})$ (5)
A value close to 1 signifies substantial visual change. For the auditory modality ($m=a$), the distance $d_{a}$ detects silence or low-energy periods directly from the raw audio signal $a_{t}$, often corresponding to semantic breaks:
$d_{a}(a_{t})=-20\log_{10}\left(\sqrt{\frac{1}{N}\sum_{i=1}^{N}a_{i}^{2}}\right)$ (6)
where $a_{i}$ are audio samples within a window and $N$ is the window size. High $d_{a}$ values indicate silence. While algorithmically distinct from neural sparse coding, this heuristic boundary detection serves the functional goal of enhancing temporal discriminability. Segment lengths are constrained between $t_{min}$ and $t_{max}$ via merging or splitting based on these detected boundaries. This segmentation process defines the temporal windows $[t_{start,i},t_{end,i}]$ for subsequent detailed encoding and memory storage.
4.2.2 Perceptual Encoding
Building upon the temporal segmentation which delineates the continuous stream into segments $i$ defined by boundaries $[t_{start,i},t_{end,i}]$, the Perceptual Encoding stage processes the raw multimodal input $x_{t}$ within each segment. Its goal is transforming the input into rich, multimodal representations suitable for memory storage and recall, mirroring the entorhinal cortex’s role in relaying sensory information to the hippocampus using powerful pre-trained models. For any given time step $t$ within a segment $i$, the system applies a tripartite encoding strategy. The visual pathway ($v$) processes $x_{t}^{v}$ using a vision-language model for descriptions $\mathcal{T}_{v}(x_{t}^{v})$. The auditory pathway ($a$) processes $x_{t}^{a}$ via speech recognition for transcriptions $\mathcal{T}_{a}(x_{t}^{a})$. The cross-modal pathway ($c$) uses a joint embedding model on $x_{t}^{v}$ and $x_{t}^{a}$ for a unified feature vector $\mathcal{E}_{c}(x_{t}^{v},x_{t}^{a})$, with optional normalization $\phi_{c}$. The collective instantaneous output is the multimodal trace $\mathcal{M}_{t}=\{\mathcal{T}_{v}(x_{t}^{v}),\mathcal{T}_{a}(x_{t}^{a}),\mathcal{E}_{c}(x_{t}^{v},x_{t}^{a})\}$.
These traces $\mathcal{M}_{t}$ are aggregated over the duration of segment $i$ (for $t$ from $t_{start,i}$ to $t_{end,i}$) to form segment-level representations: the sequence of cross-modal embeddings $\mathbf{E}_{i}=\{\mathcal{E}_{c}(x_{k}^{v},x_{k}^{a})\mid k\in[t_{start,i},t_{end,i}]\}$ and the textual descriptions $\mathbf{T}_{i}=\{\mathcal{T}_{v}(x_{k}^{v}),\mathcal{T}_{a}(x_{k}^{a})\mid k\in[t_{start,i},t_{end,i}]\}$. This aggregated content is then stored within a ShortTermMemory object $m_{i}$, which also holds pointers $\mathbf{C}_{i}$ to the raw audiovisual snippets and segment timestamps $t_{i},s_{i}$:
$m_{i}=\{\mathbf{E}_{i},\mathbf{T}_{i},\mathbf{C}_{i},t_{i},s_{i}\}$ (7)
This object $m_{i}$, representing the processed segment, serves as the short-term memory for the subsequent stage.
4.2.3 Memory Consolidation
Inspired by biological memory consolidation processes that stabilize and optimize memory traces [rolls2013quantitative], this stage aims to reduce redundancy in the sequence of ShortTermMemory objects $\{m_{i}\}$ generated by the Perceptual Encoding stage. Storing highly similar consecutive segments $m_{i},m_{i+1}$ can be inefficient. We implement a filtering mechanism based on semantic similarity in the shared cross-modal embedding space. For each segment $m_{i}$, we derive a representative cross-modal embedding $\mathbf{v}_{i}$ by averaging the sequence of embeddings $\mathbf{E}_{i}$ stored within $m_{i}$. We then maintain a set $K$ of indices corresponding to the retained, non-redundant segments. A segment $i$ is added to $K$ only if its representative feature vector $\mathbf{v}_{i}$ is sufficiently dissimilar from the representative vectors $\mathbf{v}_{j}$ of already retained segments $j$ ($j\in K,j<i$), assessed via a cosine similarity threshold $\gamma$:
$K=\{i\mid\forall j\in K,j<i\implies\cos(\mathbf{v}_{i},\mathbf{v}_{j})<\gamma\}$ (8)
Segments $i\notin K$ (i.e., those where $\cos(\mathbf{v}_{i},\mathbf{v}_{j})\geq\gamma$ for some $j\in K$) are considered redundant and discarded or merged. This sparsification reduces storage and computational load during retrieval while preserving segments that introduce significant new information, functionally analogous to interference reduction mechanisms in biological memory. The output of this stage is the sequence of consolidated ShortTermMemory objects $\{m_{k}\mid k\in K\}$.
4.2.4 Semantic Replay
Following consolidation, HippoMM implements a crucial stage inspired by neuroscientific concepts of memory replay and abstraction [eichenbaum2017prefrontal], serving as the core mechanism for transforming detailed, consolidated short-term memories into efficient, semantic long-term representations. This Semantic Replay process operates on the sequence of consolidated ShortTermMemory objects $\{m_{k}\mid k\in K\}$. Functionally, it “replays” the multimodal essence of each detailed segment $m_{k}$ through an LLM, denoted $\phi_{\text{LLM}}$, to extract and synthesize its core semantic meaning or “gist”, analogous to how hippocampal replay during sleep consolidates daytime experiences into integrated semantic knowledge in biological systems.
To generate this abstract representation, the LLM integrates key multimodal information derived from the corresponding detailed segment $m_{k}$. Specifically, the LLM processes structured multimodal context derived from $m_{k}$, including key visual information $\mathbf{I}_{\text{visual}}(m_{k})$ (extracted from visual descriptions in $\mathbf{T}_{k}$ and potentially keyframes referenced by $\mathbf{C}_{k}$) and relevant auditory information $\mathbf{A}_{\text{audio}}(m_{k})$ (extracted from transcriptions in $\mathbf{T}_{k}$), formatted to facilitate semantic summarization:
$\mathbf{S}_{\theta_{k}}=\phi_{\text{LLM}}(\text{FormattedContext}(m_{k}))$ (9)
The resulting textual summary $\mathbf{S}_{\theta_{k}}$ encapsulates the high-level semantic content distilled from the detailed segment $m_{k}$.
This semantic summary forms the centerpiece of a ThetaEvent object $\theta_{k}$, which constitutes the abstracted long-term memory representation corresponding to the detailed short-term memory $m_{k}$:
$\theta_{k}=\{\mathbf{v}_{k},\mathbf{T}_{\theta_{k}},\mathbf{I}_{\theta_{k}},\mathbf{A}_{\theta_{k}},\mathbf{S}_{\theta_{k}}\}$ (10)
Each ThetaEvent object $\theta_{k}$, representing an abstracted long-term memory representation, bundles the representative cross-modal embedding $\mathbf{v}_{k}$ (identical to the one computed during consolidation), the segment’s defining temporal information $\mathbf{T}_{\theta_{k}}$ (e.g., start/end timestamps derived from $m_{k}$’s $t_{k},s_{k}$), references to key visual ($\mathbf{I}_{\theta_{k}}$) and auditory ($\mathbf{A}_{\theta_{k}}$) context used during summarization, and the core semantic summary $\mathbf{S}_{\theta_{k}}$. This crucial temporal information $\mathbf{T}_{\theta_{k}}$ provides grounding and serves as the linkage mechanism back to the detailed consolidated ShortTermMemory object $m_{k}$, enabling transitions between abstract and detailed recall. The collection of these ThetaEvent objects $\{\theta_{k}\}$ constitutes HippoMM’s abstracted long-term memory store, facilitating efficient gist-based semantic search and reasoning while complementing the detailed episodic information retained in the consolidated short-term store $\{m_{k}\}$. This explicit transformation from detailed traces ($m_{k}$) to abstract summaries ($\theta_{k}$) operationalizes a key aspect of biological short-term to long-term memory systems, enabling more scalable and flexible memory querying. The ThetaEvent name draws inspiration from hippocampal theta rhythms’ role in temporal organization and memory sequencing [buzsaki2002theta], without explicit oscillatory modeling.
Table 1: Performance comparison of HippoMM against existing methods and ablation variants. PT: Processing Time; ART: Average Response Time; A+V: Cross-modal (Audio+Visual) accuracy; A: Audio-only accuracy; V: Visual-only accuracy; S: Semantic understanding accuracy. Best results are in bold, second best are underlined. HippoMM significantly outperforms prior methods in all modality-specific tasks while maintaining efficient processing and response times. Ablation studies demonstrate the importance of each component: Detailed Recall (DR), Fast Retrieval (FR), and Adaptive Reasoning (AR).
Method PT $\downarrow$ ART $\downarrow$ Modality Performance Avg. Acc. $\uparrow$
A+V $\uparrow$ A $\uparrow$ V $\uparrow$ S $\uparrow$
Prior Methods
NotebookLM 28.40% 23.20% 28.00% 26.80% 26.60%
Video RAG 9.46h 112.5s 63.6% 67.2% 41.2% 84.8% 64.2%
Ablation Studies
HippoMM w/o DR, AR 5.09h 4.14s 66.8% 73.2% 60.4% 90.0% 72.6%
HippoMM w/o FR, AR 5.09h 27.3s 72.0% 80.0% 66.8% 83.2% 75.5%
HippoMM w/o AR 5.09h 11.2s 68.8% 80.8% 65.6% 92.0% 76.8%
HippoMM (Ours) 5.09h 20.4s 70.8% 81.6% 66.8% 93.6% 78.2%
4.3 Memory Retrieval
4.3.1 Hierarchical Retrieval Architecture with Query Analysis
Memory retrieval commences with an analysis of the user query $q$ to discern its nature and inform subsequent retrieval steps. A query classifier, $\mathcal{Q}_{\text{type}}(q)$, determines the query’s primary modality focus (e.g., Visual $V$, Auditory $A$, Cross-modal $V\times A$) or identifies it as targeting high-level semantic/gist information. To balance retrieval efficiency and depth, HippoMM employs a hierarchical strategy that prioritizes the faster pathway. It always begins by attempting Fast Retrieval ($\Phi_{\text{fast}}$), operating on the set of abstract semantic summaries =$\mathcal{E}=\{\mathbf{S}_{\theta_{k}}\}$ derived from the ThetaEvent objects (Figure 2b, upper path). This pathway is optimized for rapidly answering semantic or gist-level queries and directly generates a candidate answer along with a confidence score. Only if the confidence score of the result from $\Phi_{\text{fast}}$ falls below a predefined threshold $\tau$, does the system escalate to the more resource-intensive Detailed Recall ($\Psi_{\text{detailed}}$). This second stage operates on the richer, consolidated ShortTermMemory objects =$\mathcal{M}=\{m_{k}\mid k\in K\}$ (Figure 2b, lower path). Crucially, should this escalation occur, the specific mechanisms employed within Detailed Recall are explicitly guided by the query type determined during the initial analysis ($\mathcal{Q}_{\text{type}}(q)$), ensuring an appropriate search strategy is used. The overall retrieval process $\mathcal{R}(q)$ follows this confidence-gated hierarchy:
$\mathcal{R}(q)=\begin{cases}\Phi_{\text{fast}}(\mathcal{E},q)&\text{if }\text{conf}(\Phi_{\text{fast}})>\tau\\ \Psi_{\text{detailed}}(q,\mathcal{M},\mathcal{Q}_{\text{type}}(q))&\text{otherwise}\end{cases}$ (11)
The hyperparameter $\tau$ solely mediates the switch between these answer-generating pathways, while the pre-determined query type directs the execution approach within $\Psi_{\text{detailed}}$ when invoked.
4.3.2 Detailed Recall Pathways
The Detailed Recall process ($\Psi_{\text{detailed}}$), invoked when Fast Retrieval yields insufficient confidence (Sec. 4.3.1), utilizes the query’s modality focus (Visual $V$, Auditory $A$, or Cross-modal $V\times A$) — identified by the classifier $\mathcal{Q}_{\text{type}}(q)$ during the initial query analysis phase — to select the appropriate search strategy within the consolidated ShortTermMemory objects $\{m_{k}\}$. The pathway chosen depends on this predetermined query type:
For single-modality queries (targeting primarily visual or auditory information), search within the relevant content of $\{m_{k}\}$ proceeds in two potential paths. First, if the query targets perceptual similarity and can be represented as an embedding $q_{embed}$, embedding-based retrieval (Feature Search) attempts a direct match using feature similarity against the representative embeddings $\{\mathbf{v}_{k}\}$ or potentially the finer-grained embedding sequences $\{\mathbf{E}_{k}\}$ stored within $\{m_{k}\}$. If this feature-based search fails, is inapplicable, or yields low confidence, text-based retrieval (Semantic Search) is performed using the aggregated textual descriptions $\mathbf{T}_{k}$ (containing captions, transcriptions) within each $m_{k}$.
For cross-modal queries (requiring association between modalities), retrieval leverages temporal co-occurrence as a proxy for episodic binding. The process unfolds in three steps: First, the system uses the pre-computed query embedding $q_{embed}$ representing the query cue. It identifies the top-$k$ segments $m_{k}$ whose representative cross-modal embedding $\mathbf{v}_{k}$ is most similar to $q_{embed}$, finding segments related to the query cue modality:
$\mathbf{S}_{\text{query}}$ $=\text{TopK}(\text{sim}(q_{embed},\{\mathbf{v}_{k}\mid m_{k}\in\mathcal{M}\}),k)$ (12)
Second, it defines expanded temporal windows $\mathbf{W}$ around the timestamps $(t_{s,k},t_{e,k})$ of these top-$k$ seed segments, using a temporal buffer $\delta$ to capture adjacent context:
$\mathbf{W}$ =$=\{[t_{s,k}-\delta,t_{e,k}+\delta]\mid k\in\mathbf{S}_{\text{query}}\}$ (13)
Third, the system retrieves information from the target modality (complementary to the query cue) by accessing segments $m_{j}$ that temporally overlap with any window in $\mathbf{W}$. This leverages temporal proximity for cross-modal association, retrieving, for example, visual details ($\mathbf{T}_{j}^{v},\mathbf{C}_{j}^{v}$) associated with an auditory cue within the same time frame:
$\mathbf{S}_{\text{target}}$ =$=\{\text{TargetInfo}(m_{j},\text{modality}_{\text{target}})\mid(t_{s,j},t_{e,j})\text{ overlaps with }\mathbf{W}\}$ (14)
where $\text{TargetInfo}(m_{j},\text{modality}_{\text{target}})$ denotes extracting the specified target modality’s content (e.g., text, embeddings, pointers) from segment $m_{j}$, and the condition $(t_{s,j},t_{e,j})\text{ overlaps with }\mathbf{W}$ is shorthand for overlapping with any interval $[t_{1},t_{2}]\in\mathbf{W}$. The hyperparameters $k$ and $\delta$ control the scope and temporal tolerance of this cross-modal retrieval mechanism.
4.3.3 Adaptive Reasoning for Final Answer Synthesis
Regardless of the retrieval path (Fast Retrieval or Detailed Recall), the collected evidence $r_{\text{retrieved}}$ must be synthesized into a final answer $a$. This is performed by the Adaptive Reasoning module $\Gamma$ (Figure 2b, “Adaptive Reasoning”), corresponding to the “AR” component (Table 1):
$a=\mathcal{R}_{\text{final}}(q)=\Gamma(q,r_{\text{retrieved}},\mathcal{E}_{\text{context}})$ (15)
$\Gamma$ is implemented using an LLM that receives query $q$, retrieved evidence $r_{\text{retrieved}}$, and context $\mathcal{E}_{\text{context}}$ with essential visual and auditory information from $r_{\text{retrieved}}$. The LLM synthesizes responses by analyzing evidence against the query, prioritizing relevance, resolving inconsistencies, and acknowledging limitations. This enables comprehensive responses that go beyond simple aggregation of the retrieved content.
5 Experiments
5.1 Implementation Details
Our system, HippoMM, was implemented on a NVIDIA L40s GPU. For Memory Formation, adaptive temporal segmentation processed video frames at 1-10 fps, identifying boundaries using SSIM and audio energy thresholds to create 5-10s segments. Perceptual Encoding generated ShortTermMemory objects ($m_{i}$, Eq. 7) containing 1024-dim ImageBind [girdhar2023imagebind] embeddings ($\mathcal{E}_{c}$), Whisper [radford2023robust] transcriptions ($\mathcal{T}_{a}$), Qwen2.5-VL [Qwen2.5-VL] visual descriptions ($\mathcal{T}_{v}$), and raw data pointers ($\mathbf{C}_{i}$). Memory Consolidation filtered segments based on cosine similarity between average ImageBind embeddings ($\mathbf{v}_{i}$) using threshold $\gamma$. Semantic Replay employed Qwen2.5-VL as $\phi_{\text{LLM}}$ to generate textual summaries ($\mathbf{S}_{\theta_{k}}$) stored in persistent ThetaEvent objects ($\theta_{k}$). For Memory Retrieval, GPT-4o [hurst2024gpt] served as the query classifier ($\mathcal{Q}_{\text{type}}$), directing queries initially to Fast Retrieval ($\Phi_{\text{fast}}$) over ThetaEvent summaries. If confidence was below $\tau=0.75$, Detailed Recall ($\Psi_{\text{detailed}}$) was invoked, using ImageBind embeddings for feature search and GPT-4o for semantic search over descriptions ($\mathbf{T}_{k}$). GPT-4o was also used as $\Gamma$ for final answer synthesis. Additional hyperparameters ($\tau_{v}$, $\tau_{a}$, $k$, $\delta$) are detailed in the Appendix.
Figure 3: Visualization of consolidated ThetaEvent embedding space and query retrieval via t-SNE projection. (Left) Embeddings for visual features, auditory features, captions, and transcriptions in one event. The semantic summary ($\mathbf{S}_{\theta}$, green star) is central. (Right) Text query embeddings (‘dorm’, ‘football’ - stars) retrieve closest caption/transcription embeddings (crosses), linking queries to specific multimodal segments (corresponding frames and text shown).
5.2 Main Results
We evaluated HippoMM on the HippoVlog benchmark (introduced in Section 3) designed for long-form audiovisual understanding, comparing it against NotebookLM [^2] https://notebooklm.google.com/ and Video RAG [ren2024videorag]. These baselines were selected as they represent the only existing systems capable of processing both long-form video and audio content. As summarized in Table 1, HippoMM achieves a state-of-the-art average accuracy of 78.2%, significantly outperforming the next-best method (Video RAG) by 14.0%. This advantage is particularly evident in visual tasks (66.8% vs. Video RAG’s 41.2%), which we attribute to our adaptive memory consolidation preserving salient details more effectively than standard chunking. Furthermore, HippoMM demonstrates strong cross-modal reasoning (70.8% A+V accuracy), confirming its ability to capture inter-sensory temporal relationships [teyler2007hippocampal]. Note that PT and ART metrics are not reported for NotebookLM, as it operates as a web-based service potentially using pre-indexed knowledge, making direct comparison infeasible. In terms of efficiency, HippoMM’s adaptive sampling reduces preprocessing time (PT) by 46.2% compared to Video RAG (5.09h vs 9.46h). Critically, HippoMM maintains a balanced average response time (ART) of 20.4s. In contrast, Video RAG exhibits a much higher ART (112.5s), largely because its retrieval mechanism fails to locate relevant segments in 55.4% of cases, defaulting to costly detailed processing across the entire long video. HippoMM’s hierarchical retrieval avoids this pitfall, balancing responsiveness and accuracy.
5.3 Ablation Studies
To assess the contribution of individual components, we performed ablation studies by selectively removing Detailed Recall (DR), Fast Retrieval (FR), and Adaptive Reasoning (AR), with results in Table 1 (bottom rows). Removing detailed, frame-level analysis (variant ‘w/o DR, AR’) achieved the fastest ART (4.14s) but at the cost of a >6% drop in visual accuracy (to 60.4%), underscoring DR’s role in fine-grained visual understanding. Conversely, removing fast, summary-based access (variant ‘w/o FR, AR’) slightly improved cross-modal accuracy (to 72.0%) but drastically increased ART to 27.3s, confirming FR’s effectiveness for rapid gist retrieval. Finally, removing only the adaptive reasoning module (‘w/o AR’) resulted in a noticeable drop in overall accuracy (to 76.8%) and semantic understanding (-1.6%), highlighting AR’s function in refining and reconciling information from different retrieval paths. These ablations demonstrate that the synergy between rapid summary access (FR), detailed segment analysis (DR), and intelligent result integration (AR) is crucial for HippoMM’s robust performance. The experimental results collectively validate our central hypothesis: translating principles inspired by human memory yields a computational framework with significant advantages in accuracy and efficiency for complex, long-form audiovisual understanding tasks.
5.4 Analysis of HippoMM Mechanisms
5.4.1 Visualizing Consolidated Event Representations and Retrieval
HippoMM employs memory consolidation and semantic replay to transform sequences of short-term memories into coherent, semantically rich ThetaEvent representations. Figure 3 provides a visualization of one such ThetaEvent, projected into a 2D space using t-SNE based on ImageBind embeddings. The left panel shows embeddings from different constituents of the event: visual features (purple squares), auditory features (orange diamonds), derived captions (blue markers), and audio transcriptions (pink circles). Notably, the embedding of the generated summary ($\mathbf{S}_{\theta}$, green star) occupies a central position relative to the various modality-specific feature clusters, suggesting its effectiveness in integrating and balancing information across different perceptual inputs. The right panel demonstrates the utility of this representation for query-driven retrieval. Text queries, such as “dorm” (yellow star query embedding) and “football” (purple star query embedding), are projected into the same space. The system identifies the nearest neighbor embeddings within the detailed ShortTermMemory features (marked by yellow and purple crosses respectively). For the “dorm” query, the system successfully retrieves a semantically relevant transcription about “stay” which, while not explicitly containing the word “dorm,” correctly identifies the residential setting depicted in the corresponding visual frame. For the “football” query, it precisely retrieves a transcription explicitly mentioning the “first football game” and correctly identifies the associated football field visual scene. This illustrates how the consolidated ThetaEvent structure supports both abstract summarization and effective retrieval of specific, grounded details from the original multimodal stream.
5.4.2 Hierarchical Retrieval Dynamics
Our hierarchical retrieval (Section 4.3) aims to balance speed and accuracy by first attempting fast, summary-based retrieval ($\Phi_{\text{fast}}$) before resorting to detailed segment search ($\Psi_{\text{detailed}}$). We analyzed the usage and performance of these pathways across the HippoVlog benchmark queries, visualizing the results in Figure 4. This figure displays individual query outcomes plotted against their response times, differentiating pathways by color and correctness by marker style (filled for correct, open for incorrect). The analysis revealed that fast retrieval successfully handled 69.1% of queries (visible by the density of points below the 30s threshold), demonstrating the effectiveness of the semantic summaries in ThetaEvent objects for rapid information access. Detailed retrieval was primarily invoked for queries requiring fine-grained perceptual information, constituting the remaining 30.9% of cases (points at or above 30s). The dashed line marks the 30s threshold. As shown in the figure annotations, the two pathways exhibit significant differences in processing efficiency, with fast retrieval achieving an average response time of 6.5s compared to detailed retrieval’s 51.6s. This clear bimodal distribution validates our hierarchical design, confirming the system effectively routes queries to the appropriate pathway based on information needs.
Figure 4: Retrieval pathway analysis with x-axis showing response time. Blue markers ($<$30s) and purple markers ($\geq$30s) represent fast and detailed pathways respectively; filled markers indicate correct outcomes, open markers incorrect. Dashed line marks the 30s threshold.
5.5 Generalization Across Audiovisual Tasks
To assess HippoMM’s generalization capabilities and benchmark its performance against contemporary methods, we established a multifaceted evaluation protocol as detailed in Table 2. This protocol integrates metrics from two complementary domains to thoroughly evaluate multimodal understanding. First, we assess event representation fidelity through Content Relevance, Semantic Similarity, and Action Alignment metrics, leveraging ground-truth action scripts from the Charades dataset [sigurdsson2016hollywood] (a detailed collection of annotated home activity videos) with evaluations supported by GPT-4o [hurst2024gpt]. Second, we probe the long-duration temporal understanding of models via the Needle QA accuracy metric, derived from the challenging Needle-in-a-Haystack task in the MLVU benchmark [MLVU], which involves locating precise information within lengthy video sequences ($\geq$ 5 min) through 415 targeted queries. The comparison includes leading multimodal systems such as Qwen2.5-Omni [xu2025qwen2] and VideoLlama2 [damonlpsg2024videollama2], alongside GPT-4o serving as a ceiling reference. Evaluation constraints include GPT-4o processing only subsampled frames (0.5 fps), and Qwen2.5-Omni’s inability to complete the Needle QA task due to memory constraints.
Results clearly highlight HippoMM’s strong generalization capabilities. Individual metrics are reported along with a consolidated Overall Score, computed as the arithmetic mean after linearly normalizing the original 1–5 scales to a 0–100% range (note that Qwen2.5-Omni’s score excludes Needle QA). As presented in Table 2, HippoMM achieves the highest Overall Score (66.9%) among directly comparable systems. Crucially, HippoMM demonstrates balanced strengths, excelling in detailed visual metrics such as Action Alignment (top score of 3.14), while significantly outperforming all other evaluated methods on the demanding long-video Needle QA task (73.1%). Moreover, HippoMM closely approximates GPT-4o’s ceiling performance, underscoring its robust visual comprehension. Collectively, these findings emphasize HippoMM’s exceptional ability to generalize across complex and diverse multimodal contexts.
Table 2: Comparison of mean evaluation scores. R: Relevance, S: Similarity, A: Alignment (Scores 1-5). NQA: Needle QA (%), Overall: Overall Score (%). GPT-4o is included as a ceiling reference for visual understanding performance. Higher values indicate better performance. Best results among directly comparable systems are in bold, second best are underlined.
Method R S A NQA (%) Overall (%)
Qwen2.5-Omni 3.39 3.14 2.76 62.0
VideoLlama2 2.32 2.17 1.59 49.9 42.9
HippoMM 3.28 3.30 3.14 73.1 66.9
GPT-4o 3.62 3.44 3.02 64.8 66.6
6 Conclusion
We introduced HippoMM, a hippocampus-inspired architecture that translates neurobiological principles into computational mechanisms for multimodal memory. Our short-to-long term memory formation process enables efficient consolidation of perceptual details into semantic abstractions. While HippoMM outperforms existing methods on audiovisual benchmarks, it struggles with ambiguous event boundaries and complex cross-modal associations under noisy conditions. The dual-pathway retrieval system effectively balances efficiency and detail but requires further refinement to handle truly open-domain queries. This work demonstrates both the promise and limitations of neuroinspired approaches, revealing that despite progress in biomimetic design, substantial gaps remain between computational implementations and the robust, flexible memory capabilities observed in biological systems.
Appendix
Appendix A HippoVlog Construction Details
The HippoVlog dataset was created through a systematic pipeline designed to capture rich multimodal information from daily vlogs and generate challenging memory-based questions. The process follows five main stages: video selection, multimodal data sampling, ground truth generation, question generation, and curation.
A.1 Video Selection and Preprocessing
  • Source Selection: 25 daily vlogs from YouTube (Creative Commons licensed), with average duration of 15.3 minutes, totaling approximately 682 minutes of content.
  • Quality Criteria: Videos were chosen based on clear audio and visual quality, natural episodic structure typical of daily life, and diverse activities (cooking, shopping, travel).
  • Format Standardization: All source videos were preprocessed to a standard format (720p resolution at 30fps) with normalized audio levels.
A.2 Multimodal Data Sampling
To create manageable and representative data points from the full vlogs, we implemented a robust sampling strategy for both audio and video modalities:
  • Audio Segment Extraction:
    • 10 non-overlapping audio segments (each 10 seconds long) were randomly selected per video.
    • Sampling avoided the initial and final 5% of videos to exclude typical intros/outros.
    • Silent segments were discarded and resampled.
  • Video Segment Extraction:
    • Corresponding 10-second video segments were extracted for each valid audio segment.
    • Checks prevented significant overlap between sampled segments.
    • Metadata about the sampled segments was preserved for future reference.
A.3 Segment-Level Ground Truth Generation
Detailed descriptions combining audio and visual information were generated for each sampled 10-second video segment:
  • Model Used: Qwen2.5-Omni-7B.
  • Process: The model processed each video segment to generate comprehensive captions.
  • Structured Output: For each segment, a structured caption containing:
    • Detailed audiovisual description
    • List of distinct visual elements
    • Transcription of speech or description of other audio elements
    • All content maintained anonymity (using “the person”, “the speaker”)
A.4 Question Generation
Multiple-choice questions were generated across four categories, with visual questions specifically based on individual frames extracted from video segments:
  • Models Used: GPT-4o.
  • Input: For visual questions, the model received frames and corresponding video context.
  • Design Principles: Questions were designed to:
    • Focus on specific, distinctive details
    • Be potentially answerable if the video had been previously seen
    • Maintain anonymity
    • Have exactly one correct answer among four plausible options
A.5 Question Curation and Finalization
The automatically generated questions underwent further curation:
  • Initial Pool: Approximately 1,500 candidate questions generated across all types.
  • Quality Control:
    • Manual validation by human.
    • Questions that were ambiguous, subjective, or had multiple plausible answers were removed or refined.
  • Final Dataset: 1,000 high-quality questions, balanced across four categories (250 per category).
  • Answer Generation: Each question includes one correct answer and three plausible but verifiably incorrect distractors, balanced for length and style.
A.6 Question Categories and Examples
Each of the four question categories in HippoVlog targets a specific memory function:
A.6.1 Cross-Modal Binding (TV×$\times$×A)
Questions requiring integration of information across visual and auditory modalities.
Example: Cross-Modal Binding Question: “When the main character is sitting in the car wearing a black jacket and white scarf, what does she audibly express?” Options: A: She says she feels excited about the drive. B: She mentions being scared and says it feels weird. C: She talks about the weather being too cold. D: She expresses confidence about her driving skills. Explanation: In the video, as the main character sits in the car wearing a black jacket and white scarf, she audibly expresses her nervousness by saying ‘okay, hey, I’m scared, oh my gosh this feels weird.’
A.6.2 Auditory-Focused Retrieval (TA)
Questions focusing on auditory information, including speech content, background sounds, and music.
Example: Auditory-Focused Retrieval Question: “What did the main character mention about the purpose of the meeting?” Options: A: To finalize a business deal B: To discuss the upcoming project C: To plan a surprise party D: To talk about a vacation Explanation: The main character specifically mentions the ‘upcoming project’ as the purpose of the meeting, indicating a discussion centered around future plans and tasks.
A.6.3 Visual-Focused Retrieval (TV)
Questions targeting visual details that require careful observation of scene elements.
Example: Visual-Focused Retrieval Question: “What distinctive pattern is visible on the roof of the building?” Options: A: Checkerboard B: Chevron C: Striped D: Polka dot Explanation: The roof of the building in the image has a distinctive chevron pattern.
Example: Visual-Focused Retrieval Question: “What is the main character wearing on her head?” Options: A: A baseball cap B: A knitted hat C: A colorful headscarf D: A headband Explanation: The main character is wearing a colorful headscarf, as seen in the image.
A.6.4 Semantic/Temporal Reasoning (TS)
Questions requiring integration of information across temporal segments and abstract understanding.
Example: Semantic/Temporal Reasoning Question: “What was one significant challenge the main character faced during the experiment?” Options: A: Finding time to eat breakfast B: Dealing with brain fog C: Avoiding social media D: Lack of exercise options Explanation: The main character starts the day with a homemade green juice made from various fruits and vegetables.
Appendix B HippoMM Implementation Details
This appendix provides the specific implementation choices, models, and parameters used in the HippoMM system, complementing the conceptual description in Section 4.
System Configuration
All experiments reported were conducted on a single NVIDIA L40s GPU.
B.1 Memory Formation Implementation
Temporal Pattern Separation (Sec. 4.2.1)
The process segments continuous video streams based on perceptual changes:
  • Input Sampling: Raw video frames are adaptively sampled at a rate between 1 and 10 frames per second (fps). Raw audio is sampled at 16 kHz PCM.
  • Visual Change Detection: A boundary is triggered ($\mathcal{S}_{t}=1$) if the visual dissimilarity between consecutive frames, measured as $d_{v}(F_{t},F_{t-1})=1-\text{SSIM}(F_{t},F_{t-1})$ (Eq. 5), exceeds a threshold $\tau_{v}=0.65$.
  • Auditory Change Detection: A boundary is triggered ($\mathcal{S}_{t}=1$) if the negative log energy of the raw audio signal, $d_{a}(a_{t})$ (Eq. 6), exceeds a threshold $\tau_{a}=40$. This corresponds to the signal amplitude dropping below a threshold (approx. -40 dB), indicating potential silence or a significant break.
  • Segment Duration: The resulting segments are constrained to be between $t_{min}=5$ seconds and $t_{max}=10$ seconds long. Shorter segments following boundary detection are merged with preceding ones, and longer intervals between boundaries are split to adhere to this constraint, defining the temporal windows $[t_{start,i},t_{end,i}]$.
Perceptual Encoding (Sec. 4.2.2)
Each segment $i$ (defined by $[t_{start,i},t_{end,i}]$) is processed to extract multimodal features:
  • Cross-Modal Embeddings: We use ImageBind (v1.0) as the joint embedding model ($\mathcal{E}_{c}$) to generate sequences of shared 1024-dimensional embeddings ($\mathbf{E}_{i}$) from visual frames and audio snippets within the segment.
  • Audio Transcription: Whisper (v3 medium) is employed as the speech recognition model ($\mathcal{T}_{a}$) to obtain time-aligned transcriptions from the segment’s audio.
  • Visual Description: Qwen2.5-VL is used as the vision-language model ($\mathcal{T}_{v}$) to generate textual descriptions of the visual content (maximum length 200 tokens).
  • ShortTermMemory Object ($m_{i}$): These processed elements are aggregated and stored in a ShortTermMemory object $m_{i}$, as defined in Eq. 7. It encapsulates: pointers to the raw audiovisual snippets ($\mathbf{C}_{i}$), the sequence of ImageBind embeddings ($\mathbf{E}_{i}$), the aggregated Whisper transcriptions and Qwen-VL visual descriptions ($\mathbf{T}_{i}$), segment start/end timestamps ($t_{i},s_{i}$), and related metadata.
Memory Consolidation (Sec. 4.2.3)
Redundancy is reduced by filtering similar consecutive segments:
  • Similarity Calculation: A representative cross-modal embedding ($\mathbf{v}_{i}$) is calculated for each segment $m_{i}$ by averaging its sequence of ImageBind embeddings ($\mathbf{E}_{i}$).
  • Filtering Criterion: A segment $m_{i}$ is retained in the consolidated set $K$ only if the cosine similarity between its representative embedding $\mathbf{v}_{i}$ and the representative embedding $\mathbf{v}_{j}$ of the most recent previously retained segment $m_{j}$ (where $j\in K,j<i$) is below a predefined threshold $\gamma=0.85$, as per Eq. 8. Segments not meeting this criterion are discarded (or merged). The output is the sequence of consolidated objects $\{m_{k}\mid k\in K\}$.
Semantic Replay (Sec. 4.2.4)
Abstracted representations are generated for long-term storage:
  • Summarization Model: The Qwen2.5-VL large vision-language model ($\phi_{\text{LLM}}$) processes formatted multimodal context derived from the consolidated segment $m_{k}$ (including key visual information $\mathbf{I}_{\text{visual}}(m_{k})$ and auditory information $\mathbf{A}_{\text{audio}}(m_{k})$). It generates a concise textual summary ($\mathbf{S}_{\theta_{k}}$) using a temperature of 0.7 and a maximum length of 300 tokens (see Table 3 for prompt details), as per Eq. 9.
  • ThetaEvent Object ($\theta_{k}$): This object constitutes the long-term memory entry, storing the elements defined in Eq. 10: the representative cross-modal embedding ($\mathbf{v}_{k}$), the generated textual summary ($\mathbf{S}_{\theta_{k}}$), references to the key visual ($\mathbf{I}_{\theta_{k}}$) and auditory ($\mathbf{A}_{\theta_{k}}$) context used, and the segment’s temporal information ($\mathbf{T}_{\theta_{k}}$) which links back to the detailed ShortTermMemory object $m_{k}$. The collection $\{\theta_{k}\}$ forms the abstract long-term memory store.
B.2 Memory Retrieval Implementation
Query Analysis (Sec. 4.3.1)
User queries are initially processed to determine the retrieval strategy:
  • Classifier Model: GPT-4o is used as the query classifier ($\mathcal{Q}_{\text{type}}$) to classify the incoming query $q$ based on its primary modality focus (Visual, Auditory, Cross-modal) or if it targets high-level semantic/gist information (Summary). This uses a 3-shot prompt with examples (see Table 4).
Hierarchical Retrieval (Sec. 4.3.1)
Retrieval follows a confidence-gated two-pathway approach (Eq. 11):
  • Fast Retrieval ($\Phi_{\text{fast}}$): Retrieval is first attempted using the abstract textual summaries ($\mathbf{S}_{\theta_{k}}$) stored in the collection of ThetaEvent objects (=$\mathcal{E}=\{\theta_{k}\}$). Confidence scores are determined by internal LLM calibration (Qwen2.5-VL, see Table 5) on a scale of 0-1.
  • Detailed Recall ($\Psi_{\text{detailed}}$): If the confidence score from Fast Retrieval falls below a threshold $\tau=0.75$, retrieval escalates to the Detailed Recall pathway. This pathway searches the consolidated ShortTermMemory objects (=$\mathcal{M}=\{m_{k}\mid k\in K\}$), guided by the query type determined by $\mathcal{Q}_{\text{type}}(q)$.
Detailed Recall Pathways (Sec. 4.3.2)
The specific operations within Detailed Recall ($\Psi_{\text{detailed}}$) are guided by the classified query type:
  • Query Embedding: The query $q$ is converted into an ImageBind embedding ($q_{embed}$) for use in similarity-based search.
  • Similarity Search (Feature/Semantic): For single-modality or cue-based retrieval (Visual or Auditory query type), cosine similarity is used to find the top-$k$ ShortTermMemory segments ($m_{k}$) whose representative embeddings ($\mathbf{v}_{k}$) are most similar to $q_{embed}$ (Feature Search), with $k=5$. Text-based search over the aggregated descriptions/transcriptions ($\mathbf{T}_{k}$) serves as a fallback or alternative (Semantic Search). See Tables 6 and 7 for related selection prompts.
  • Cross-Modal Association: For Cross-modal query types, retrieval uses temporal co-occurrence. First, the top-$k$ seed segments $\mathbf{S}_{\text{query}}$ most similar to $q_{embed}$ (based on $\mathbf{v}_{k}$) are identified (Eq. 12, with $k=5$). Second, expanded temporal windows ($\mathbf{W}$) are defined around these seeds using a temporal buffer parameter $\delta=2$ seconds (Eq. 13). Third, target modality information ($\mathbf{S}_{\text{target}}$) is extracted from any segment $m_{j}$ whose time interval overlaps with any window in $\mathbf{W}$ (Eq. 14).
  • Fine-grained Analysis: If necessary for answering the query, specific content within the retrieved ShortTermMemory segments is analyzed further using Qwen2.5-VL (for visual/multimodal questions) or Whisper (for detailed audio content examination). This often occurs during the final synthesis stage.
Adaptive Reasoning (Sec. 4.3.3)
The final answer is synthesized from the retrieved information:
  • Synthesis Model: Qwen2.5-VL is typically used as the reasoning module ($\Gamma$) to generate the final answer ($a$). It takes the original query ($q$), the retrieved evidence ($r_{\text{retrieved}}$), and relevant extracted context ($\mathcal{E}_{\text{context}}$) as input (Eq. 15, see Table 9 for prompt). For complex reasoning tasks requiring deeper integration or handling potential conflicts (e.g., comparing Fast Retrieval and Detailed Recall outputs), GPT-4o may be used with a 4-shot prompting strategy and temperature of 0.2 (see Table 8 for a related prompt).
Appendix C Additional Experimental Details
C.1 Ablation Studies Parameter Settings
Acronyms refer to components described in Sec. 4: FR (Fast Retrieval, $\Phi_{\text{fast}}$), DR (Detailed Recall, $\Psi_{\text{detailed}}$), AR (Adaptive Reasoning, $\Gamma$).
  • HippoMM w/o DR, AR: Uses Fast Retrieval ($\Phi_{\text{fast}}$) only, with direct answer generation from summaries (no Adaptive Reasoning $\Gamma$).
  • HippoMM w/o FR, AR: Uses Detailed Recall ($\Psi_{\text{detailed}}$) only, with direct answer generation from retrieved details (no Adaptive Reasoning $\Gamma$).
  • HippoMM w/o AR: Uses both retrieval paths ($\Phi_{\text{fast}}$, $\Psi_{\text{detailed}}$) as per the hierarchical logic, but the final output is based on direct extraction or simple aggregation rather than the full synthesis step ($\Gamma$).
C.2 Performance Metrics
  • Processing Time (PT): Total time to execute the Memory Formation phase (Sec. 4.2) on raw videos.
  • Average Response Time (ART): Average time to execute the Memory Retrieval phase (Sec. 4.3) and generate an answer $a$ for a query $q$.
  • Accuracy Metrics: Based on strict binary match with human-annotated ground truth answers.
C.3 Generalization Evaluation
  • Charades Dataset: 100 randomly selected video clips.
  • MLVU Needle-in-a-Haystack: 415 queries across 150 long-form videos (5+ minutes each).
  • GPT-4o Evaluation: Used as judge for Content Relevance, Semantic Similarity, and Action Alignment with a 5-point Likert scale.
  • Overall Score Calculation: Arithmetic mean after linearly normalizing 1-5 scales to 0-100%.
C.4 Additional Implementation Parameters
  • Key Frame Filtering Threshold: Within Perceptual Encoding, a visual similarity threshold (SSIM-based, value 0.3) is used to filter visually redundant frames before description generation to reduce processing load. (Note: This is distinct from the consolidation threshold $\gamma$).
  • Visual Similarity Fallback Threshold (Detailed Recall): When performing Feature Search in Detailed Recall for visual queries, if the max similarity score is below 0.4, the system may fall back to Semantic Search using text captions.
  • Direct Answer Confidence Threshold ($\tau$): The threshold used in Hierarchical Retrieval (Eq. 11) to switch from Fast Retrieval to Detailed Recall is $\tau=0.75$.
  • Context Length (LLM): A maximum context length of 120,000 tokens is used for GPT-4o when invoked for complex reasoning (Adaptive Reasoning $\Gamma$).
  • Multiple Worker Processing: Multiprocessing with a pool size equal to the number of available CPU cores (up to a maximum of 8) is used for parallel frame processing during Perceptual Encoding.
Table 3: Semantic Replay Summarization Prompt ($\phi_{\text{LLM}}$)
Task Description Please provide a concise one sentence summary of this event based on the video frames descriptions and audio transcription. What is happening in this event? Input Format Image descriptions: [Concatenated frame captions from key visual context $\mathbf{I}_{\text{visual}}(m_{k})$] Audio transcription: [Concatenated transcribed text from key auditory context $\mathbf{A}_{\text{audio}}(m_{k})$] Output Requirements One sentence summary ($\mathbf{S}_{\theta_{k}}$) capturing the integrated multimodal essence of the event Focus on key actions, subjects, and setting Maintain factual accuracy based on provided descriptions Avoid introducing information not supported by the input
Table 4: Query Type Classification Prompt ($\mathcal{Q}_{\text{type}}$)
Task Description Classify this question into one of these categories based on what type of information is needed to answer it: Classification Categories VIDEO - Questions specifically about visual elements, appearances, or actions that need frame-by-frame analysis, e.g., “what is the main character holding?” (Corresponds to Visual $V$) AUDIO - Questions about sounds, speech, or audio content that need audio analysis, e.g., “what does the main character say/mention?” (Corresponds to Auditory $A$) VIDEO+AUDIO - Questions that require associating both visual and audio information, e.g., “what is the main character doing while saying/mentioning something?” (Corresponds to Cross-modal $V\times A$) SUMMARY - Questions that focus on the overall content or gist of the video, e.g., “what is the main character doing?” (Corresponds to Semantic/Gist) Guidelines If question is about visual details, appearances, or actions, classify as VIDEO If question is about sounds, speech, or audio content, classify as AUDIO If question requires both visual and audio information, classify as VIDEO+AUDIO If question is about the overall content of the video, classify as SUMMARY Input Question: [user query $q$] Output Format Return ONLY one of these exact words: VIDEO, AUDIO, VIDEO+AUDIO, SUMMARY
Table 5: Fast Retrieval Attempt Prompt (Part of $\Phi_{\text{fast}}$)
Task Description Given the following question and relevant event summaries ($\mathbf{S}_{\theta_{k}}$ from $\mathcal{E}$), analyze whether the question can be answered directly or needs specific analysis (triggering Detailed Recall $\Psi_{\text{detailed}}$). Input Question: [user query $q$] Event Summaries: [List of relevant event summaries $\mathbf{S}_{\theta_{k}}$ retrieved based on $q$] Guidelines for analysis General questions about overall video content (likely classified as SUMMARY) should be attempted directly from summaries. Questions about specific visual details (VIDEO), sounds/speech (AUDIO), or combined elements (VIDEO+AUDIO) often require specific analysis beyond summaries. Estimate confidence based on whether summaries contain sufficient information. Output Format Must be in one of these two structures: 1. If answerable from summaries:
ANSWER: <your detailed answer based ONLY on summaries>
CONFIDENCE: <score between 0.0-1.0, e.g., 0.7 or higher if confident>
2. If requiring specific analysis (summaries insufficient):
ANSWER: NONE
CONFIDENCE: <low score, e.g., 0.0 or < 0.7>
(Note: This confidence score is compared against threshold $\tau$)
Table 6: Visual Segment Selection Prompt (Part of $\Psi_{\text{detailed}}$ Semantic Search)
Task Description Given a question (potentially classified as VIDEO or VIDEO+AUDIO), descriptions from consolidated memory segments ($\mathbf{T}_{k}$), and an optional element to search for, identify at most 5 relevant frame indices (within segments $m_{k}$) for answering the question. Input Question: [user query $q$] Element to search for: [search query derived from $q$] Frame descriptions (from relevant $\mathbf{T}_{k}$): [Numbered list of frame descriptions/captions] Instructions Return ONLY numbers separated by commas (e.g., “0,3,5,8,12”) corresponding to indices within the provided list. Return at most 5 indices. Do not include any other text, explanations, or spaces. If fewer than 5 frames are relevant, return fewer indices. Example good responses “0,3,5,8,12” “1,4,7” Output Format Your response (numbers only, comma-separated):
Table 7: Audio Segment Selection Prompt (Part of $\Psi_{\text{detailed}}$ Semantic Search)
Task Description Given a question (potentially classified as AUDIO or VIDEO+AUDIO) and transcriptions with timestamps from consolidated memory segments ($\mathbf{T}_{k}$), identify the most relevant time frames where the answer might be found. Input Question: [user query $q$] Transcriptions (with timestamps from relevant $\mathbf{T}_{k}$): [List of timestamped transcriptions] Instructions Analyze the transcriptions and identify segments most likely to contain the answer. Return a JSON array of time frames (relative to the source video). Each frame defines a start and end time.
    {"start": START_TIME, "end": END_TIME},
    {"start": START_TIME, "end": END_TIME}
Return at most 5 time frames. Include a small buffer around identified time frames (e.g., $\pm$2 seconds, related to $\delta$). If no relevant segments found, return “[]”.
Output Format Your response (valid JSON only):
Table 8: Adaptive Reasoning Prompt (Optional step within $\Gamma$)
Task Description You potentially have two candidate answers to the same question $q$. One derived from Fast Retrieval ($\Phi_{\text{fast}}$, using summaries $\mathcal{E}$) and another from Detailed Recall ($\Psi_{\text{detailed}}$, using detailed segments $\mathcal{M}$). Compare them and provide a final, reconciled answer $a$. Input Question: [user query $q$] Answer from summaries (if $\Phi_{\text{fast}}$ produced one): [direct_answer] Confidence from summaries: [direct_confidence] Answer from detailed analysis (if $\Psi_{\text{detailed}}$ was run): [detailed_answer] Supporting Context ($\mathcal{E}_{\text{context}}$): Sample frame captions: [Sample relevant frame captions from $r_{\text{retrieved}}$] Sample transcriptions: [Sample relevant transcriptions from $r_{\text{retrieved}}$] Instructions Compare both answers for consistency, detail, and plausibility given the context. If they agree, prefer the more detailed/confident answer. If they disagree, evaluate which answer is better supported by the specific context (captions, transcriptions). Prioritize the answer derived from detailed analysis ($\Psi_{\text{detailed}}$) if confidence in the summary-based answer ($\Phi_{\text{fast}}$) was low or if the detailed analysis provides clearly contradictory evidence. Synthesize the best possible final answer based on this comparison. For multiple choice questions, select only one final answer option. Estimate the confidence in the final reconciled answer. Output Format
ANSWER: <reconciled final answer a>
CONFIDENCE: <final confidence score between 0.0-1.0>
REASONING: <brief explanation for the final answer choice>
Table 9: Final Answer Synthesis Prompt ($\Gamma$)
Task Description Based on the following retrieved context ($r_{\text{retrieved}}$) and extracted evidence ($\mathcal{E}_{\text{context}}$), please synthesize a final answer ($a$) to the question ($q$). Input Question: [user query $q$] Overall Video Context (e.g., relevant summaries $\mathbf{S}_{\theta_{k}}$ if available): [List of relevant high-level context] Relevant Retrieved Content (details from $\Phi_{\text{fast}}$ or $\Psi_{\text{detailed}}$): [List of retrieved content descriptions, e.g., frame captions, transcriptions, details from $m_{k}$] Instructions Please provide a clear and specific answer based on the integration of the overall context and the specific retrieved content. If the information is insufficient or contradictory, state that or provide the best possible reasoned answer based on the evidence. Output should be one letter if the original question was multiple choice. Output Format Answer: <Synthesized answer $a$>

### v2

@@
 Abstract
-Comprehending extended audiovisual experiences remains a fundamental challenge for computational systems.
-Current approaches struggle with temporal integration and cross-modal associations that humans accomplish effortlessly through hippocampal-cortical networks.
-We introduce HippoMM, a biologically-inspired architecture that transforms hippocampal mechanisms into computational advantages for multimodal understanding.
-HippoMM implements three key innovations: (i) hippocampus-inspired pattern separation and completion specifically designed for continuous audiovisual streams, (ii) short-to-long term memory consolidation that transforms perceptual details into semantic abstractions, and (iii) cross-modal associative retrieval pathways enabling modality-crossing queries.
-Unlike existing retrieval systems with static indexing schemes, HippoMM dynamically forms integrated episodic representations through adaptive temporal segmentation and dual-process memory encoding.
-Evaluations on our challenging HippoVlog benchmark demonstrate that HippoMM significantly outperforms state-of-the-art approaches (78.2% vs.
-64.2% accuracy) while providing substantially faster response times (20.4s vs.
-112.5s).
-Our results demonstrate that translating neuroscientific memory principles into computational architectures provides a promising foundation for next-generation multimodal understanding systems.
-The code and benchmark dataset are publicly available at https://github.com/linyueqian/HippoMM.
+Comprehending extended audiovisual experiences remains challenging for computational systems, particularly temporal integration and cross-modal associations fundamental to human episodic memory.
+We introduce HippoMM, a computational cognitive architecture that maps hippocampal mechanisms to solve these challenges.
+Rather than relying on scaling or architectural sophistication, HippoMM implements three integrated components: (i) Episodic Segmentation detects audiovisual input changes to split videos into discrete episodes, mirroring dentate gyrus pattern separation; (ii) Memory Consolidation compresses episodes into summaries with key features preserved, analogous to hippocampal memory formation; and (iii) Hierarchical Memory Retrieval first searches semantic summaries, then escalates via temporal window expansion around seed segments for cross-modal queries, mimicking CA3 pattern completion.
+These components jointly create an integrated system exceeding the sum of its parts.
+On our HippoVlog benchmark testing associative memory, HippoMM achieves state-of-the-art 78.2% accuracy while operating 5x faster than retrieval-augmented baselines.
+Our results demonstrate that cognitive architectures provide blueprints for next-generation multimodal understanding.
 1 Introduction
-Imagine hearing a familiar melody through an open window, instantly recalling not merely the music but the complete concert experience: the interplay of stage lights, the energy of the crowd, and even the aroma of summer rain.
-This remarkable cognitive capability, known as multimodal pattern completion, involves reconstructing integrated experiences from partial, cross-modal stimuli and remains a significant frontier in computational modeling.
-Although current methods proficiently process static multimodal data [radford2021learning, girdhar2023imagebind] or retrieve relevant segments from extensive video corpora [ren2024videorag], they struggle fundamentally with long-form multimodal comprehension.
-Specifically, these approaches exhibit limitations in three key dimensions: segmenting continuous audiovisual streams into coherent episodic units, reconstructing rich experiences by integrating cross-modal and temporal information (particularly from partial cues), and flexibly retrieving information across modalities—such as recalling visual details from auditory cues.
-In contrast, the human hippocampus robustly achieves these capabilities.
-As illustrated conceptually in Figure 1 (top), the hippocampus integrates multisensory inputs (e.g., visual and auditory streams) through structures including the entorhinal cortex, organizes continuous experiences into discrete episodes [baldassano2017discovering], employs pattern separation mechanisms typically attributed to the Dentate Gyrus (DG), and performs pattern completion associated primarily with CA3 circuits to store and recall distinct memory traces [yassa2011pattern, rolls2013quantitative].
-These processes ultimately yield cohesive episodic representations, mediated by hippocampal circuitry including the CA1 region [eichenbaum2014time].
-Figure 1: Conceptual overview of hippocampal versus HippoMM multimodal processing.
-(Top) Biological hippocampus integrates visual and auditory modalities through the entorhinal cortex and associated circuits (DG, CA3, CA1) to form and recall episodic memories.
-(Bottom) Proposed HippoMM architecture processes multimodal inputs (video, audio, cross-modal) inspired by hippocampal principles for episodic memory formation and cue-driven retrieval.
-Despite this clear biological blueprint, existing computational memory systems often mimic only isolated aspects.
-They typically operate unimodally [gutierrez2024hipporag, graves2016hybrid] or function as passive retrieval systems [lewis2020retrieval, ren2024videorag], falling short of replicating the active, reconstructive, and crossmodal associative nature of hippocampal memory, particularly for extended audiovisual data streams.
-To address these limitations, we introduce HippoMM, a novel architecture designed to computationally implement functional principles inspired by hippocampal mechanisms for long form multimodal content understanding, as conceptualized in Figure 1 (bottom).
-Processing video, audio, and jointly learned multimodal representations, HippoMM aims to emulate the brain’s ability to form and query rich, integrated multimodal memories.
-While prior work has drawn inspiration from the hippocampus for unimodal tasks or specific memory functions, HippoMM represents, to our knowledge, the first architecture to systematically adapt and integrate a suite of hippocampal memory mechanisms to address the unique challenges of long-form audiovisual understanding.
-Our contributions are as follows:
-  1. 1.
-    Hippocampal Dynamics for Long AV Streams: We operationalize dentate gyrus-inspired pattern separation and CA3-like pattern completion specifically for continuous audio-visual data.
-This allows HippoMM to segment lengthy streams into distinct episodes and reconstruct coherent multimodal experiences even from fragmented cues, moving beyond passive retrieval.
-  2. 2.
-    Short-to-Long Term Memory Consolidation: We model the transition from detailed perceptual traces (short-term) to abstract semantic representations (long-term) through a neuro-inspired consolidation and semantic replay process.
-This contrasts with models that use fixed chunking or lack mechanisms for evolving representations over time, enabling more efficient and meaningful long-term storage.
-  3. 3.
-    Cross-Modal Associative Querying: HippoMM introduces retrieval pathways that explicitly support querying information in one modality using cues from another (e.g., finding what was seen when a specific sound occurred, or vice-versa).
-This capability, absent in prior multimodal systems that often disregard audio richness or lack such cross-modal retrieval tests, mimics a key aspect of human associative recall.
-Unlike existing approaches that treat modalities in isolation, our system maintains the temporal co-occurrence relationships essential for reconstructing coherent multimodal experiences from partial cues.
-Evaluated on our proposed HippoVlog benchmark, comprising long-form vlogs with challenging cross-modal and unimodal question-answering tasks, HippoMM achieves state-of-the-art accuracy (78.2% average accuracy, significantly outperforming VideoRAG [ren2024videorag]), particularly excelling in cross-modal and visual reasoning tasks.
-Ablation studies underscore the critical roles of biologically inspired components: removing detailed recall mechanisms reduces visual accuracy by 6%, and eliminating adaptive reasoning results in a 1.6% decline in semantic understanding.
-The remainder of the paper is structured as follows: Section 2 discusses related research on hippocampal mechanisms and multimodal understanding; Section 3 formalizes the multimodal memory-augmented retrieval task and introduces the HippoVlog benchmark; Section 4 describes the HippoMM architecture in detail; Section 5 presents comprehensive experimental results; and finally, we conclude with discussions on implications and future directions in Section 6.
-Our work demonstrates that integrating hippocampal principles can advance computational understanding of complex, continuous multimodal experiences, offering a robust foundation toward developing systems with human-like memory and reasoning capabilities.
+Despite recent progress, current multimodal models still cannot perform long-form audiovisual understanding accurately and efficiently.
+Specifically, they (i) must retain entire continuous content spanning several hours as the context [42], (ii) cannot reconstruct complete experiences when given partial sensory cues [17], and (iii) cannot form abstract memories that transfer across contexts [35].
+Such failures prevent machines from seamlessly understanding real-world experiences the way humans naturally do: we effortlessly (i) segment raw, continuous content into meaningful episodes [40], (ii) recall entire episodes from a single sound or image [33], and (iii) extract lasting knowledge from fleeting perceptions [23].
+Figure 1: Conceptual Overview of Hippocampal vs.
+HippoMM Processing.
+The biological hippocampus (top) and our HippoMM architecture (bottom) both integrate continuous sensory streams to form and recall episodic memories.
+While prior works improve multimodal systems by scaling up or introducing sophisticated architectures [26, 9], recent memory-augmented approaches [14] show explicit memory mechanisms are crucial.
+We posit that current systems lack effective cognitive mechanisms for episodic memory.
+The biological hippocampus provides a cognitive architecture that solves these challenges through three core mechanisms: segmentation of continuous experience into discrete episodes, associative reconstruction from fragmentary inputs, and consolidation of perceptual details into semantic knowledge [38, 29].
+Figure 1 illustrates the parallel between biological hippocampal processing and our computational approach.
+By translating these principles into algorithmic implementations, we design an accurate and efficient multimodal system.
+To this end, in this work we propose HippoMM, which implements our insight through direct functional mapping of hippocampal computations.
+Following dentate gyrus (DG) pattern separation [38], our Episodic Segmentation employs content-adaptive boundaries to parse inputs into episodic codes.
+Emulating hippocampal consolidation processes [16], Memory Consolidation progressively transforms perceptual traces into semantic representations, culminating in ThetaEvent memories.
+Mimicking CA3 autoassociation [29], Hierarchical Memory Retrieval enables pattern completion from partial cues and cross-modal queries.
+This integrated architecture amplifies each component through their interactions, achieving performance gains that validate the cognitive architectural approach.
+Another contribution of our work is HippoVlog, a benchmark designed to test cross-modal associative recall in long-form video understanding.
+Unlike existing benchmarks that focus on sequential comprehension [22] or simple retrieval tasks [39], HippoVlog evaluates episodic memory functions through partial cue reconstruction and semantic consolidation questions.
+Built from 25 daily vlogs, it provides 1,000 manually-validated questions across four memory function categories.
+In summary, our contributions are:
+  • We design HippoMM, which implements hippocampal memory functions through episodic segmentation at perceptual boundaries, cross-modal feature binding in consolidated memory units, and hierarchical retrieval that balances semantic summaries with detailed recall.
+  • We introduce HippoVlog, a benchmark for cross-modal associative memory with questions requiring pattern completion from partial cues, addressing the gap in existing benchmarks that only test sequential comprehension.
+  • We demonstrate that cognitive architectures provide principled solutions: HippoMM achieves 78.2% accuracy and speedup over retrieval baselines, validating biological mechanisms as effective computational designs.
 2 Related Work
 2.1 Hippocampal Memory Mechanisms
-Neuroscientific research highlights the hippocampus’s crucial ability to form, organize, and retrieve episodic memories.
-Foundational work established key principles including pattern separation through sparse coding in the dentate gyrus [yassa2011pattern], autoassociative pattern completion in CA3 [rolls2013quantitative], and theta-gamma oscillations for temporal coding [lisman2009prediction].
-Recent findings reveal how the hippocampus segments experiences at event boundaries [baldassano2017discovering, ben2013hippocampal], creating hierarchical memory structures [zacks2007event] that optimize storage and recall.
-Computational models have translated these principles to machine learning through temporal chunking [dubrow2013influence] and neural memory networks [kumaran2016learning], but primarily in unimodal contexts.
-The hippocampus’s role as a multimodal integrator [teyler2007hippocampal, eichenbaum2014time] provides critical insights for artificial memory systems.
-Neurobiological studies demonstrate its capacity to bind cross-modal features while maintaining modality-specific details [quiroga2009explicit, mormann2008latency], a capability existing computational models lack.
-Critically, effectively integrating continuous, cross-modal audiovisual streams remains underdeveloped, even in recent memory architectures [whittington2020tolman], presenting an important research direction.
-2.2 Audio-Visual Multimodal Understanding
-Modern multimodal systems leverage joint embedding spaces [radford2021learning, girdhar2023imagebind] to align audio-visual modalities, while video-language models [alayrac2022flamingo, zhang2023video] employ temporal attention for sequential processing.
-Though effective for static retrieval, these methods lack mechanisms for temporal memory organization critical for long-form event understanding.
-Recent audio-visual models [shu2023audio, guo2025aligned] enhance modality integration through specialized tokens or multi-scale adapters, yet struggle with extended temporal contexts and cross-modal associative memory.
-Retrieval-augmented approaches [ren2024videorag, yuan2025memory] address long-context challenges through graph-based grounding and cognitive memory steps, but remain biologically uninspired.
-Cross-modal retrieval frameworks [lin2024mm] mitigate modality bias yet lack temporal organization, while efficiency-focused methods [lin2025speechprune, liu2025keyframe] prune redundant elements without enabling progressive retrieval pathways.
+The hippocampus performs three key functions: pattern separation through sparse coding in the DG [38], pattern completion via CA3 autoassociative networks [29], and event segmentation at boundaries [3].
+While computational models have adapted these principles [16], they primarily focus on unimodal contexts, overlooking the hippocampus’s role as a multimodal integrator [6].
+2.2 AudioVisual Multimodal Understanding
+Current multimodal systems employ joint embeddings [26, 9] or video-language models [1, 20] but lack mechanisms for temporal memory organization and associative recall.
+While recent models like VideoLLaMA 2 [5] and Qwen2.5-Omni [37] incorporate audiovisual processing, and voice-native benchmarks [19] expose modality-induced reasoning gaps, these efforts focus on modality alignment rather than memory formation.
+Related benchmarks such as LongVALE [8] target temporal grounding and dense captioning—predicting temporal boundaries in pre-segmented video—which is complementary to but distinct from our task of cross-modal associative recall from continuous streams.
+Critically, existing approaches cannot form episodic memories from continuous streams or perform cross-modal pattern completion, where sounds trigger visual memories or vice versa.
+This fundamental gap in hippocampal-like memory mechanisms represents a major limitation, as without them, models process only pre-segmented clips through attention [32] or fusion [24, 18, 34] without true associative recall.
 2.3 Biomimetic Memory Architectures
-Memory-augmented neural networks have evolved from early differentiable memory systems [santoro2016meta] to sophisticated architectures explicitly mirroring hippocampal-cortical interactions [whittington2020tolman].
-Hippocampus-inspired models like HippoRAG [gutierrez2024hipporag] adapt principles of hippocampal indexing theory for knowledge integration, drawing on biological mechanisms similar to the dentate gyrus’ pattern separation [yassa2011pattern] and CA3’s autoassociative recall [rolls2013quantitative] for efficient document retrieval.
-Similarly, Differentiable Neural Computers [graves2016hybrid] emulate the hippocampus’ sparse temporal coding mechanisms [lisman2009prediction].
-However, these approaches remain predominantly unimodal, neglecting the hippocampus’ critical role as a multimodal binding engine that coordinates cortical sensory streams through theta-phase coding, grid cell-inspired spatial organization, and cross-modal integration [eichenbaum2014time].
+Memory-augmented networks like the Differentiable Neural Computer [12] pioneered external memory but remain predominantly unimodal.
+Recent hippocampal-inspired architectures include HippoRAG [13] for text retrieval and MA-LMM [14] which extends memory-augmented approaches to long-term video understanding through memory banks.
+HippoMM uniquely integrates pattern separation, consolidation, and cross-modal pattern completion in a unified architecture for continuous audiovisual understanding.
+Figure 2: The HippoMM architecture.
+(a) Memory Formation: composed of Episodic Segmentation ($\mathcal{S}_{t}$) based on perceptual boundaries, Perceptual Encoding of visual and auditory inputs with cross-modal features, and Memory Consolidation that progressively transforms short-term features through similarity-based filtering ($K$) into consolidated features, feature collections, and ultimately semantic summaries to form ThetaEvent ($\theta$).
+(b) Hierarchical Memory Retrieval: implementing Query-Driven Pattern Completion through fast retrieval ($\Phi_{\text{fast}}$) and detailed recall pathways with temporal window localization ($\mathbf{W}$).
 3 Multimodal Memory-augmented Retrieval
 3.1 Task Formulation
-We formalize the Multimodal Memory-augmented Retrieval (MMR) task as the challenge of developing systems that process, understand, and reason about long-form continuous multimodal data streams $X=\{x_{t}\}$, exemplified by extended audiovisual recordings.
-Given a user query $q$ targeting visual ($q_{V}$), auditory ($q_{A}$), cross-modal ($q_{V\times A}$), or semantic ($q_{S}$) information, the system must construct and utilize appropriate memory representations to synthesize a coherent and accurate answer $a$.
-The task decomposes into three fundamental stages.
-First, a memory formation function $\mathcal{F}_{\text{mem}}$ transforms the raw input stream $X$ into a structured memory representation $M$:
-$M=\mathcal{F}_{\text{mem}}(X)$ (1)
-This memory $M=\{m_{k}\}$ comprises encoded segments $m_{k}$, each capturing multimodal features, temporal context, and semantic abstractions.
-Second, a retrieval function $\mathcal{R}_{\text{retrieval}}$ processes the query $q$ against memory $M$ to identify relevant evidence $E$:
-$E=\mathcal{R}_{\text{retrieval}}(q,M)$ (2)
-Here, $E$ typically includes a relevance-ranked subset of memory elements (i.e., $E=\{\langle m_{k},r_{k}\rangle\mid m_{k}\in M\land\text{relevant}(m_{k},q)\}$).
-Finally, an answer synthesis function $\mathcal{G}_{\text{synth}}$ generates the response $a$ by reasoning over the query $q$ and retrieved evidence $E$:
-$a=\mathcal{G}_{\text{synth}}(q,E)$ (3)
-Implementing an effective MMR system requires sophisticated realizations of $\mathcal{F}_{\text{mem}}$, $\mathcal{R}_{\text{retrieval}}$, and $\mathcal{G}_{\text{synth}}$ that handle temporal dependencies, enable cross-modal associations, and support robust reasoning, analogous to biological memory processes described in cognitive neuroscience [yassa2011pattern, eichenbaum2014time].
-3.2 HippoVlog Dataset
-To evaluate MMR systems under realistic conditions, we introduce HippoVlog.
-Existing benchmarks for multimodal understanding typically utilize short video clips, rely solely on Large Language Model (LLM)-based evaluation without ground truth answers [ren2024videorag], or lack audio elements for cross-modal reasoning within continuous, long-form videos [MLVU].
-HippoVlog addresses these gaps, comprising 25 long-form daily vlogs (682 minutes total) with naturalistic audiovisual content.
-It includes 1,000 validated multiple-choice question-answer pairs, each with four candidate answers, targeting specific memory functions:
-  • Cross-modal binding ($T_{V\times A}$): Linking visual/auditory cues (e.g., "What object appeared when the speaker mentioned hiking?").
-  • Auditory-focused retrieval ($T_{A}$): Extracting audio details (e.g., "What instrument played during the introduction?").
-  • Visual-focused retrieval ($T_{V}$): Extracting visual details (e.g., "What text was visible on the store sign?").
-  • Semantic reasoning ($T_{S}$): Integrating semantic information over time (e.g., "What are the cooking steps mentioned throughout the video?").
-Questions were created using GPT-4o [hurst2024gpt], manually validated, and annotated with timestamps.
-HippoVlog’s objective assessment format, focus on audio-visual integration, and long-form temporal dynamics make it particularly suitable for evaluating MMR systems.
-Construction details are provided in the Appendix and our dataset is publicly available [^1] https://github.com/linyueqian/HippoVlog/.
-Figure 2: The HippoMM architecture for multimodal memory.
-(a) Memory Formation: composed of Temporal Pattern Separation ($\mathcal{S}_{t}$) based on perceptual boundaries, Perceptual Encoding of visual and auditory inputs with cross-modal features, Memory Consolidation using similarity-based filtering ($K$), and Semantic Replay generating ThetaEvent representations ($\theta$).
-(b) Memory Retrieval: implementing Query-Driven Pattern Completion through fast retrieval ($\Phi_{\text{fast}}$) and detailed recall pathways with temporal window localization ($\mathbf{W}$).
-4 HippoMM System
-4.1 Overview
-We introduce HippoMM, a computational architecture designed for the complex task of understanding and reasoning about long-form multimodal data streams, typified by extended audiovisual recordings.
-The architecture draws functional inspiration from the mammalian hippocampus’s established roles in episodic memory processing [eichenbaum2017prefrontal], including cross-modal binding, temporal sequence organization, pattern separation (differentiating similar inputs), pattern completion (retrieving full memories from partial cues), and memory consolidation.
-While HippoMM does not model neural dynamics directly, it seeks to replicate key functional principles through distinct algorithmic modules.
-Our objective is to overcome the limitations of conventional multimodal systems, which often lack robust mechanisms for handling continuous temporal dynamics and associative cross-modal recall.
-HippoMM enables context-sensitive retrieval, facilitating the reconstruction of integrated multimodal experiences from potentially incomplete queries (e.g., retrieving visual context associated with a specific sound event).
-The system architecture, depicted schematically in Figure 2, comprises two primary phases: Memory Formation (Sec.
-4.2), which processes raw inputs into structured, abstracted representations in long-term memory (Figure 2a), and Memory Retrieval (Sec.
-4.3), which leverages these representations to answer user queries via hierarchical and associative search strategies (Figure 2b).
+We formalize the task of Multimodal Memory-augmented Retrieval (MMR).
+While retrieval tasks have been extensively studied across various modalities, existing formulations typically treat retrieval as matching queries against pre-indexed content.
+Our work uniquely addresses retrieval from continuous multimodal streams requiring active memory formation, which is fundamentally different from traditional settings as the system must continuously construct and maintain structured memory representations from ongoing inputs, mimicking human episodic memory.
+The objective is to develop a system that processes a long-form, continuous multimodal data stream $X=\{x_{t}\}$ and, given a query $q$, synthesizes an accurate answer $a$ through three functional stages: (i) memory formation $\mathcal{F}_{\text{mem}}$ transforms $X$ into structured representation $M$; (ii) retrieval $\mathcal{R}_{\text{retrieval}}$ identifies relevant evidence $E$ from $M$; and (iii) answer synthesis $\mathcal{G}_{\text{synth}}$ generates response $a$.
+HippoMM instantiates $\mathcal{F}_{\text{mem}}$ via content-adaptive segmentation, cross-modal encoding, and similarity-filtered consolidation; $\mathcal{R}_{\text{retrieval}}$ via GPT-4o-based confidence-gated hierarchical retrieval; and $\mathcal{G}_{\text{synth}}$ via LLM-based adaptive reasoning—each grounded in hippocampal principles but not the only valid instantiation.
+3.2 The HippoVlog Benchmark
+A significant challenge in evaluating MMR systems is the lack of benchmarks designed to test for cross-modal associative recall.
+Recent benchmarks have advanced long-form video understanding significantly: MLVU [43] evaluates multi-task comprehension across diverse genres, ALLVB [31] unifies nine tasks with 2-hour videos, and Video-MME [7] provides comprehensive multimodal evaluation.
+However, these benchmarks fundamentally test comprehension of presented information rather than memory formation and associative recall.
+Specifically, existing benchmarks evaluate whether models can answer questions about explicitly shown content, but they do not test: (i) the ability to form episodic memories from continuous experience that enable later reconstruction from partial cues, (ii) cross-modal pattern completion where a sound triggers recall of associated visual content or vice versa, and (iii) the consolidation of perceptual details into transferable semantic knowledge.
+These capabilities, central to human episodic memory, require fundamentally different evaluation paradigms than sequential comprehension.
+To address this gap, we introduce HippoVlog, a benchmark specifically designed to evaluate hippocampal-like memory functions.
+Built from 25 long-form daily vlogs totaling 682 minutes, it provides 1,000 questions across four memory function categories: Cross-modal binding ($T_{V\times A}$), Auditory-focused retrieval ($T_{A}$), Visual-focused retrieval ($T_{V}$), and Semantic reasoning ($T_{S}$).
+The quality of the benchmark is ensured through rigorous manual validation, achieving a high inter-annotator agreement (Cohen’s $\kappa=0.975$).
+Unlike comprehension-focused benchmarks, HippoVlog’s questions probe whether systems can reconstruct complete multimodal experiences from fragmentary cues—the hallmark of episodic memory.
+Construction details are provided in the Appendix.
+Our code is publicly available.
+[^1] Code: https://github.com/linyueqian/HippoMM
+4 Methodology
+We implement the fundamental cognitive architecture of the hippocampus, translating its three core computations of pattern separation (DG), pattern completion (CA3), and memory consolidation (CA1) into algorithms for multimodal understanding.
+This hippocampal-inspired architecture addresses precisely those challenges that confound current models: segmenting continuous experience into discrete episodes, reconstructing complete memories from partial cues, and abstracting semantic knowledge from perceptual details.
+Our methodology implements these biological principles through a two-phase process: a Memory Formation phase that mirrors the hippocampal encoding pathway from entorhinal cortex through DG-CA3-CA1, and a Hierarchical Memory Retrieval phase that implements CA3 autoassociation and CA1 semantic recall for query-driven reasoning.
 4.1 Memory Formation
-4.2.1 Temporal Pattern Separation
-Continuous experience is often perceived and remembered as discrete events [baldassano2017discovering].
-To computationally mimic the hippocampus’s ability to segment experience (a function often linked to the dentate gyrus [yassa2011pattern]), HippoMM first employs content-adaptive temporal segmentation as a functional approximation.
-This stage analyzes the incoming raw audiovisual stream to identify boundaries corresponding to significant perceptual shifts, partitioning the stream into potential event segments.
-This contrasts with fixed-duration chunking, aiming instead for more semantically coherent and efficiently processed units.
-A segment boundary is indicated at time $t$ by $\mathcal{S}_{t}=1$ if a significant change is detected in either the visual or auditory stream, using computationally efficient metrics applied directly to the input:
-$\mathcal{S}_{t}=\bigvee_{m\in\{v,a\}}\text{1}\left[d_{m}(\text{Input}_{m}(t),\text{Input}_{m}(t-1))>\tau_{m}\right]$ (4)
-where $d_{m}$ is a modality-specific distance function comparing relevant inputs for modality $m$ at times $t$ and $t-1$, $\tau_{m}$ is a predefined threshold, $\text{1}[\cdot]$ is the indicator function, and $\bigvee$ denotes logical OR.
-Specifically, for the visual modality ($m=v$), we define the distance $d_{v}$ based on the Structural Similarity Index Measure (SSIM) between consecutive raw visual frames $F_{t},F_{t-1}$:
-$d_{v}(F_{t},F_{t-1})=1-\text{SSIM}(F_{t},F_{t-1})$ (5)
-A value close to 1 signifies substantial visual change.
-For the auditory modality ($m=a$), the distance $d_{a}$ detects silence or low-energy periods directly from the raw audio signal $a_{t}$, often corresponding to semantic breaks:
-$d_{a}(a_{t})=-20\log_{10}\left(\sqrt{\frac{1}{N}\sum_{i=1}^{N}a_{i}^{2}}\right)$ (6)
-where $a_{i}$ are audio samples within a window and $N$ is the window size.
-High $d_{a}$ values indicate silence.
-While algorithmically distinct from neural sparse coding, this heuristic boundary detection serves the functional goal of enhancing temporal discriminability.
-Segment lengths are constrained between $t_{min}$ and $t_{max}$ via merging or splitting based on these detected boundaries.
-This segmentation process defines the temporal windows $[t_{start,i},t_{end,i}]$ for subsequent detailed encoding and memory storage.
-4.2.2 Perceptual Encoding
-Building upon the temporal segmentation which delineates the continuous stream into segments $i$ defined by boundaries $[t_{start,i},t_{end,i}]$, the Perceptual Encoding stage processes the raw multimodal input $x_{t}$ within each segment.
-Its goal is transforming the input into rich, multimodal representations suitable for memory storage and recall, mirroring the entorhinal cortex’s role in relaying sensory information to the hippocampus using powerful pre-trained models.
-For any given time step $t$ within a segment $i$, the system applies a tripartite encoding strategy.
-The visual pathway ($v$) processes $x_{t}^{v}$ using a vision-language model for descriptions $\mathcal{T}_{v}(x_{t}^{v})$.
-The auditory pathway ($a$) processes $x_{t}^{a}$ via speech recognition for transcriptions $\mathcal{T}_{a}(x_{t}^{a})$.
-The cross-modal pathway ($c$) uses a joint embedding model on $x_{t}^{v}$ and $x_{t}^{a}$ for a unified feature vector $\mathcal{E}_{c}(x_{t}^{v},x_{t}^{a})$, with optional normalization $\phi_{c}$.
-The collective instantaneous output is the multimodal trace $\mathcal{M}_{t}=\{\mathcal{T}_{v}(x_{t}^{v}),\mathcal{T}_{a}(x_{t}^{a}),\mathcal{E}_{c}(x_{t}^{v},x_{t}^{a})\}$.
-These traces $\mathcal{M}_{t}$ are aggregated over the duration of segment $i$ (for $t$ from $t_{start,i}$ to $t_{end,i}$) to form segment-level representations: the sequence of cross-modal embeddings $\mathbf{E}_{i}=\{\mathcal{E}_{c}(x_{k}^{v},x_{k}^{a})\mid k\in[t_{start,i},t_{end,i}]\}$ and the textual descriptions $\mathbf{T}_{i}=\{\mathcal{T}_{v}(x_{k}^{v}),\mathcal{T}_{a}(x_{k}^{a})\mid k\in[t_{start,i},t_{end,i}]\}$.
-This aggregated content is then stored within a ShortTermMemory object $m_{i}$, which also holds pointers $\mathbf{C}_{i}$ to the raw audiovisual snippets and segment timestamps $t_{i},s_{i}$:
-$m_{i}=\{\mathbf{E}_{i},\mathbf{T}_{i},\mathbf{C}_{i},t_{i},s_{i}\}$ (7)
-This object $m_{i}$, representing the processed segment, serves as the short-term memory for the subsequent stage.
-4.2.3 Memory Consolidation
-Inspired by biological memory consolidation processes that stabilize and optimize memory traces [rolls2013quantitative], this stage aims to reduce redundancy in the sequence of ShortTermMemory objects $\{m_{i}\}$ generated by the Perceptual Encoding stage.
-Storing highly similar consecutive segments $m_{i},m_{i+1}$ can be inefficient.
-We implement a filtering mechanism based on semantic similarity in the shared cross-modal embedding space.
-For each segment $m_{i}$, we derive a representative cross-modal embedding $\mathbf{v}_{i}$ by averaging the sequence of embeddings $\mathbf{E}_{i}$ stored within $m_{i}$.
-We then maintain a set $K$ of indices corresponding to the retained, non-redundant segments.
-A segment $i$ is added to $K$ only if its representative feature vector $\mathbf{v}_{i}$ is sufficiently dissimilar from the representative vectors $\mathbf{v}_{j}$ of already retained segments $j$ ($j\in K,j<i$), assessed via a cosine similarity threshold $\gamma$:
-$K=\{i\mid\forall j\in K,j<i\implies\cos(\mathbf{v}_{i},\mathbf{v}_{j})<\gamma\}$ (8)
-Segments $i\notin K$ (i.e., those where $\cos(\mathbf{v}_{i},\mathbf{v}_{j})\geq\gamma$ for some $j\in K$) are considered redundant and discarded or merged.
-This sparsification reduces storage and computational load during retrieval while preserving segments that introduce significant new information, functionally analogous to interference reduction mechanisms in biological memory.
-The output of this stage is the sequence of consolidated ShortTermMemory objects $\{m_{k}\mid k\in K\}$.
-4.2.4 Semantic Replay
-Following consolidation, HippoMM implements a crucial stage inspired by neuroscientific concepts of memory replay and abstraction [eichenbaum2017prefrontal], serving as the core mechanism for transforming detailed, consolidated short-term memories into efficient, semantic long-term representations.
-This Semantic Replay process operates on the sequence of consolidated ShortTermMemory objects $\{m_{k}\mid k\in K\}$.
-Functionally, it “replays” the multimodal essence of each detailed segment $m_{k}$ through an LLM, denoted $\phi_{\text{LLM}}$, to extract and synthesize its core semantic meaning or “gist”, analogous to how hippocampal replay during sleep consolidates daytime experiences into integrated semantic knowledge in biological systems.
-To generate this abstract representation, the LLM integrates key multimodal information derived from the corresponding detailed segment $m_{k}$.
-Specifically, the LLM processes structured multimodal context derived from $m_{k}$, including key visual information $\mathbf{I}_{\text{visual}}(m_{k})$ (extracted from visual descriptions in $\mathbf{T}_{k}$ and potentially keyframes referenced by $\mathbf{C}_{k}$) and relevant auditory information $\mathbf{A}_{\text{audio}}(m_{k})$ (extracted from transcriptions in $\mathbf{T}_{k}$), formatted to facilitate semantic summarization:
-$\mathbf{S}_{\theta_{k}}=\phi_{\text{LLM}}(\text{FormattedContext}(m_{k}))$ (9)
-The resulting textual summary $\mathbf{S}_{\theta_{k}}$ encapsulates the high-level semantic content distilled from the detailed segment $m_{k}$.
-This semantic summary forms the centerpiece of a ThetaEvent object $\theta_{k}$, which constitutes the abstracted long-term memory representation corresponding to the detailed short-term memory $m_{k}$:
-$\theta_{k}=\{\mathbf{v}_{k},\mathbf{T}_{\theta_{k}},\mathbf{I}_{\theta_{k}},\mathbf{A}_{\theta_{k}},\mathbf{S}_{\theta_{k}}\}$ (10)
-Each ThetaEvent object $\theta_{k}$, representing an abstracted long-term memory representation, bundles the representative cross-modal embedding $\mathbf{v}_{k}$ (identical to the one computed during consolidation), the segment’s defining temporal information $\mathbf{T}_{\theta_{k}}$ (e.g., start/end timestamps derived from $m_{k}$’s $t_{k},s_{k}$), references to key visual ($\mathbf{I}_{\theta_{k}}$) and auditory ($\mathbf{A}_{\theta_{k}}$) context used during summarization, and the core semantic summary $\mathbf{S}_{\theta_{k}}$.
-This crucial temporal information $\mathbf{T}_{\theta_{k}}$ provides grounding and serves as the linkage mechanism back to the detailed consolidated ShortTermMemory object $m_{k}$, enabling transitions between abstract and detailed recall.
-The collection of these ThetaEvent objects $\{\theta_{k}\}$ constitutes HippoMM’s abstracted long-term memory store, facilitating efficient gist-based semantic search and reasoning while complementing the detailed episodic information retained in the consolidated short-term store $\{m_{k}\}$.
-This explicit transformation from detailed traces ($m_{k}$) to abstract summaries ($\theta_{k}$) operationalizes a key aspect of biological short-term to long-term memory systems, enabling more scalable and flexible memory querying.
-The ThetaEvent name draws inspiration from hippocampal theta rhythms’ role in temporal organization and memory sequencing [buzsaki2002theta], without explicit oscillatory modeling.
-Table 1: Performance comparison of HippoMM against existing methods and ablation variants.
+Our memory formation pipeline implements three computational primitives from the hippocampus: pattern separation (episodic segmentation), perceptual encoding, and memory consolidation (semantic abstraction).
+These stages transform continuous multimodal streams into a hierarchical memory structure with both detailed episodic traces and efficient semantic indices.
+Episodic Segmentation (Pattern Separation / Threshold-based Boundary Detection).
+Following the DG’s pattern separation function that transforms overlapping inputs into distinct neural codes, we implement content-adaptive temporal segmentation to prevent interference between similar memories.
+Fixed-window chunking arbitrarily divides videos at regular intervals, often splitting coherent events mid-action or grouping unrelated scenes together.
+To preserve semantic integrity, our architecture segments the stream by detecting perceptual discontinuities that align with natural event boundaries.
+Specifically, we trigger a boundary at time $t$ when either visual similarity drops significantly between consecutive frames or level drops below a threshold indicating silence or scene breaks:
+$\mathcal{S}_{t}=\text{1}[d_{v}(F_{t},F_{t-1})>\tau_{v}]\vee\text{1}[d_{a}(a_{t})>\tau_{a}],$ (1)
+where visual dissimilarity $d_{v}$ is measured using Structural Similarity Index Measure (SSIM) (Eq.
+2) with threshold $\tau_{v}$ implementing the sparsity constraint observed in DG neurons, and audio boundary detection $d_{a}$ (Eq.
+3) uses threshold $\tau_{a}$ chosen through empirical validation to identify pauses in speech or sound that often coincide with activity transitions.
+$d_{v}(F_{t},F_{t-1})$ $=1-\text{SSIM}(F_{t},F_{t-1}),$ (2)
+$d_{a}(a_{t})$ $=-20\log_{10}\left(\sqrt{\frac{1}{N}\sum_{i=1}^{N}a_{i}^{2}}\right).$ (3)
+These perceptually-motivated boundaries produce segments constrained between $t_{min}=5$ and $t_{max}=10$ seconds, aligning with human event segmentation timescales reported in prior work [40].
+This content-aware segmentation produces a series of coherent temporal windows, $t_{s,i},t_{e,i}]$, forming the foundational units for memory encoding.
+Perceptual Encoding.
+Mirroring how the entorhinal cortex provides multimodal input transformed through DG into CA3’s distributed representations, we employ specialized encoders that preserve both modality-specific features and cross-modal associations.
+Within each event segment, the system captures a rich, multimodal representation of the experience.
+This is not a monolithic encoding but a tripartite strategy that orchestrates specialized foundation models: a vision-language model generates dense textual descriptions ($\mathcal{T}_{v}$), a speech recognition model provides precise transcriptions ($\mathcal{T}_{a}$), and a cross-modal embedding model ($\mathcal{E}_{c}$) maps both visual frames and audio snippets into a shared feature space, enabling direct similarity comparisons across modalities.
+This distributed encoding mirrors CA3’s recurrent architecture where diverse inputs converge into unified episodic representations.
+These outputs are aggregated into a detailed episodic record, the ShortTermMemory object $m_{i}$:
+$m_{i}=\{\mathbf{E}_{i},\mathbf{T}_{i},\mathbf{C}_{i},t_{s,i},t_{e,i}\},$ (4)
+where $\mathbf{E}_{i}$ denotes the sequence of cross-modal embeddings from $\mathcal{E}_{c}$, $\mathbf{T}_{i}$ aggregates the textual outputs from both $\mathcal{T}_{v}$ (visual descriptions) and $\mathcal{T}_{a}$ (audio transcriptions), $\mathbf{C}_{i}$ contains pointers to the raw audiovisual data, and $t_{s,i},t_{e,i}$] marks the temporal boundaries of segment $i$.
+Memory Consolidation (Synaptic Consolidation / Cosine-Similarity Sparse Filtering).
+Implementing synaptic and systems consolidation (with CA3/CA1 roles), our pipeline progressively abstracts from perceptual details to semantic summaries.
+Storing every fine-grained detail from a long stream is inefficient and leads to catastrophic interference between similar events.
+Our architecture addresses this with a consolidation mechanism that actively filters for novelty.
+This process filters redundant ShortTermMemory objects, retaining only those that introduce significant new information.
+For each segment $m_{i}$, a representative embedding $\mathbf{v}_{i}$ is computed by averaging the sequence of cross-modal embeddings $\mathbf{E}_{i}$ produced by the embedding model $\mathcal{E}_{c}$.
+The segment is retained in the consolidated set $K$ only if its embedding is sufficiently dissimilar from all previously stored memories, governed by a similarity threshold $\gamma$:
+$K=\{i\mid\forall j\in K,j<i\implies\cos(\mathbf{v}_{i},\mathbf{v}_{j})<\gamma\}.$ (5)
+This threshold $\gamma$ is chosen to encourage sparse retention, analogous to CA3 sparsity where only 2-5% of neurons are active for any given memory.
+This selective filtering creates an efficient memory store that is robust to distraction and optimized for retrieval.
+The consolidated set $K$ contains indices of retained segments.
+For clarity, we use $k\in K$ to denote consolidated segment indices and $m_{k}$ for their corresponding memory objects.
+The cornerstone of our architecture’s efficiency is the creation of a high-level semantic index over the consolidated episodic records.
+This stage transforms detailed, fine-grained memories into abstract, language-based summaries, making vast stretches of time searchable at near-zero cost.
+For each consolidated memory $m_{k}$ where $k\in K$ denotes indices from the filtered set, an LLM ($\phi_{\text{LLM}}$) synthesizes its multimodal contents into a concise textual “gist,” $\mathbf{S}_{\theta_{k}}$:
+$\mathbf{S}_{\theta_{k}}=\phi_{\text{LLM}}(\text{FormattedContext}(m_{k})),$ (6)
+where $\text{FormattedContext}(\cdot)$ structures the multimodal content from $m_{k}$ (visual descriptions, transcriptions, and temporal information) into a prompt format suitable for the LLM.
+This summary becomes the centerpiece of a ThetaEvent object, our long-term memory data structure.
+The name “ThetaEvent” reflects hippocampal theta oscillations that bind experiences into coherent episodes.
+The ThetaEvent bundles the abstract summary with the representative embedding $\mathbf{v}_{k}$ and a pointer back to its detailed source memory, creating a powerful dual-representation system that bridges semantic abstraction and perceptual detail, precisely the function performed by CA1 in biological memory consolidation.
+$\theta_{k}=\{\mathbf{v}_{k},\mathbf{S}_{\theta_{k}},\mathbf{T}_{\theta_{k}},\mathbf{I}_{\theta_{k}},\mathbf{A}_{\theta_{k}}\},$ (7)
+where $\mathbf{v}_{k}$ is the representative embedding, $\mathbf{S}_{\theta_{k}}$ is the generated semantic summary, $\mathbf{T}_{\theta_{k}}$ contains temporal references linking to the source $m_{k}$, and $\mathbf{I}_{\theta_{k}}$, $\mathbf{A}_{\theta_{k}}$ store key visual and auditory context used in summary generation.
+4.2 Hierarchical Memory Retrieval (Pattern Completion / Confidence-Gated Two-Path Lookup)
+The hippocampus supports two complementary retrieval modes: rapid semantic access through CA1 and detailed pattern completion through CA3 recurrence.
+Our retrieval architecture directly implements this biological duality, resolving the critical trade-off between speed and fidelity by dynamically choosing its search strategy based on the query’s demands, facilitating complex forms of reasoning such as associative recall and cross-modal pattern completion.
+Query-Driven Pathway Selection.
+Implementing the hippocampus’s dual retrieval modes: CA1’s rapid semantic access and CA3’s detailed pattern completion, we employ confidence-gated hierarchical retrieval.
+Upon receiving a query $q$, the system first attempts Fast Retrieval ($\Phi_{\text{fast}}$), implementing CA1’s semantic pathway by searching exclusively on the abstract ThetaEvent summaries.
+This path rapidly resolves high-level and gist-based queries.
+Only if the confidence of this initial result falls below a tuned threshold $\tau$ does the system escalate to the more resource-intensive Detailed Recall ($\Psi_{\text{detailed}}$), which engages CA3’s pattern completion mechanisms.
+This confidence score is generated by the LLM’s internal calibration when attempting to answer from summaries alone, reflecting its assessment of whether sufficient information exists in the abstract representations.
+This on-demand escalation provides a “zoom-in” capability, accessing the rich, consolidated ShortTermMemory objects for evidence-grounded reasoning.
+$\mathcal{R}(q)=\begin{cases}\Phi_{\text{fast}}(\Theta,q)&\text{if }\text{conf}(\Phi_{\text{fast}})>\tau,\\ \Psi_{\text{detailed}}(q,\mathcal{M},\mathcal{Q}_{\text{type}}(q))&\text{otherwise},\end{cases}$ (8)
+where $\Theta=\{\theta_{k}\mid k\in K\}$ denotes the collection of ThetaEvent objects.
+Detailed Recall with Cross-Modal Pattern Completion.
+CA3’s recurrent connectivity enables pattern completion: reconstructing complete memories from partial cues.
+This biological computation is most powerful for cross-modal binding, where a cue in one modality triggers recall of associated information in another modality.
+When detailed recall is invoked, the system deploys a specialized search strategy tailored to the query’s modality.
+For single-modality queries, it performs either an embedding-based Feature Search or a text-based Semantic Search.
+The key innovation, however, is its mechanism for cross-modal associative recall, which uses temporal co-occurrence as a powerful proxy for episodic binding, mirroring how CA3 neurons bind coincident inputs through synaptic plasticity.
+This process begins by using the query cue to find relevant seed segments in the target memory store:
+$\mathbf{S}_{\text{query}}=\text{TopK}(\text{sim}(q_{embed},\{\mathbf{v}_{k}\mid m_{k}\in\mathcal{M}\}),k).$ (9)
+It then defines expanded temporal windows $\mathbf{W}$ around these seeds to capture the surrounding event context, implementing the temporal receptive fields of hippocampal place cells.
+$\mathbf{W}=\{[t_{s,k}-\delta,t_{e,k}+\delta]\mid k\in\mathbf{S}_{\text{query}}\}.$ (10)
+Finally, it retrieves information from the target modality by searching for any memory segments that temporally overlap with these windows, effectively reconstructing the integrated experience.
+$\mathbf{S}_{\text{target}}=\{\text{Extract}_{\text{target}}(m_{j})\mid\\ (t_{s,j},t_{e,j})\text{ overlaps with }\mathbf{W}\},$ (11)
+where $\text{Extract}_{\text{target}}(m_{j})$ extracts modality-specific content from segment $m_{j}$: visual descriptions from $\mathbf{T}_{j}$ for visual targets, or transcriptions for auditory targets.
+We denote segment $j$’s temporal boundaries as $(t_{s,j},t_{e,j})$.
+To illustrate cross-modal pattern completion, consider the query: “What was on the screen when the applause started?” The system identifies the auditory cue (“applause”) and visual target (“on the screen”), with low confidence from Fast Retrieval triggering Detailed Recall due to the specific temporal link.
+Following Eq.
+9, it finds seed segment $m_{k}$ containing the applause sound, then forms temporal window $\mathbf{W}$ around it (Eq.
+10)—implementing how hippocampal neurons encode temporally extended episodes.
+Visual information from segments overlapping $\mathbf{W}$ is retrieved (Eq.
+11) and passed to Adaptive Reasoning for synthesis.
+Table 1: Performance comparison of HippoMM against existing methods and ablation variants on HippoVlog.
 PT: Processing Time; ART: Average Response Time; A+V: Cross-modal (Audio+Visual) accuracy; A: Audio-only accuracy; V: Visual-only accuracy; S: Semantic understanding accuracy.
@@
 Prior Methods
-NotebookLM 28.40% 23.20% 28.00% 26.80% 26.60%
-Video RAG 9.46h 112.5s 63.6% 67.2% 41.2% 84.8% 64.2%
+NotebookLM [11] 28.40% 23.20% 28.00% 26.80% 26.60%
+VideoRAG [28] 9.46h 112.5s 63.6% 67.2% 41.2% 84.8% 64.2%
+Ola [21] 79.4s 72.4% 85.6% 57.6% 84.0% 74.9%
+VideoLLaMA 3 [41] 58.3s - - 70.8% 75.2% 73.0%
+Simple-Assoc-Memory (SAM) 32.0% 32.8% 33.2% 23.2% 30.3%
+Gemini-3-Flash [10] 44.0% 45.2% 40.0% 56.0% 46.3%
+GPT-5 [25] 72.0% 73.2% 45.6% 88.0% 69.7%
+Qwen2.5-Omni [37] 71.2% 69.6% 51.2% 73.6% 66.4%
 Ablation Studies
@@
 HippoMM w/o AR 5.09h 11.2s 68.8% 80.8% 65.6% 92.0% 76.8%
+HippoMM w/o DR 5.09h 6.39s 39.2% 65.6% 48.0% 92.0% 61.2%
+HippoMM w/o FR 5.09h 19.54s 72.0% 79.6% 63.6% 83.2% 74.6%
+HippoMM (Qwen2.5-14B) 5.09h 15.7s 69.3% 77.1% 56.6% 76.4% 70.8%
+HippoMM (EOR-only) 5.09h 63.2% 70.4% 60.8% 90.0% 71.1%
+HippoMM ($\gamma=0.80$) 5.09h 20.4s 69.2% 78.4% 64.0% 90.8% 75.6%
+HippoMM ($\gamma=0.90$) 5.09h 20.4s 71.6% 78.8% 64.4% 90.0% 76.2%
 HippoMM (Ours) 5.09h 20.4s 70.8% 81.6% 66.8% 93.6% 78.2%
-4.3 Memory Retrieval
-4.3.1 Hierarchical Retrieval Architecture with Query Analysis
-Memory retrieval commences with an analysis of the user query $q$ to discern its nature and inform subsequent retrieval steps.
-A query classifier, $\mathcal{Q}_{\text{type}}(q)$, determines the query’s primary modality focus (e.g., Visual $V$, Auditory $A$, Cross-modal $V\times A$) or identifies it as targeting high-level semantic/gist information.
-To balance retrieval efficiency and depth, HippoMM employs a hierarchical strategy that prioritizes the faster pathway.
-It always begins by attempting Fast Retrieval ($\Phi_{\text{fast}}$), operating on the set of abstract semantic summaries =$\mathcal{E}=\{\mathbf{S}_{\theta_{k}}\}$ derived from the ThetaEvent objects (Figure 2b, upper path).
-This pathway is optimized for rapidly answering semantic or gist-level queries and directly generates a candidate answer along with a confidence score.
-Only if the confidence score of the result from $\Phi_{\text{fast}}$ falls below a predefined threshold $\tau$, does the system escalate to the more resource-intensive Detailed Recall ($\Psi_{\text{detailed}}$).
-This second stage operates on the richer, consolidated ShortTermMemory objects =$\mathcal{M}=\{m_{k}\mid k\in K\}$ (Figure 2b, lower path).
-Crucially, should this escalation occur, the specific mechanisms employed within Detailed Recall are explicitly guided by the query type determined during the initial analysis ($\mathcal{Q}_{\text{type}}(q)$), ensuring an appropriate search strategy is used.
-The overall retrieval process $\mathcal{R}(q)$ follows this confidence-gated hierarchy:
-$\mathcal{R}(q)=\begin{cases}\Phi_{\text{fast}}(\mathcal{E},q)&\text{if }\text{conf}(\Phi_{\text{fast}})>\tau\\ \Psi_{\text{detailed}}(q,\mathcal{M},\mathcal{Q}_{\text{type}}(q))&\text{otherwise}\end{cases}$ (11)
-The hyperparameter $\tau$ solely mediates the switch between these answer-generating pathways, while the pre-determined query type directs the execution approach within $\Psi_{\text{detailed}}$ when invoked.
-4.3.2 Detailed Recall Pathways
-The Detailed Recall process ($\Psi_{\text{detailed}}$), invoked when Fast Retrieval yields insufficient confidence (Sec.
-4.3.1), utilizes the query’s modality focus (Visual $V$, Auditory $A$, or Cross-modal $V\times A$) — identified by the classifier $\mathcal{Q}_{\text{type}}(q)$ during the initial query analysis phase — to select the appropriate search strategy within the consolidated ShortTermMemory objects $\{m_{k}\}$.
-The pathway chosen depends on this predetermined query type:
-For single-modality queries (targeting primarily visual or auditory information), search within the relevant content of $\{m_{k}\}$ proceeds in two potential paths.
-First, if the query targets perceptual similarity and can be represented as an embedding $q_{embed}$, embedding-based retrieval (Feature Search) attempts a direct match using feature similarity against the representative embeddings $\{\mathbf{v}_{k}\}$ or potentially the finer-grained embedding sequences $\{\mathbf{E}_{k}\}$ stored within $\{m_{k}\}$.
-If this feature-based search fails, is inapplicable, or yields low confidence, text-based retrieval (Semantic Search) is performed using the aggregated textual descriptions $\mathbf{T}_{k}$ (containing captions, transcriptions) within each $m_{k}$.
-For cross-modal queries (requiring association between modalities), retrieval leverages temporal co-occurrence as a proxy for episodic binding.
-The process unfolds in three steps: First, the system uses the pre-computed query embedding $q_{embed}$ representing the query cue.
-It identifies the top-$k$ segments $m_{k}$ whose representative cross-modal embedding $\mathbf{v}_{k}$ is most similar to $q_{embed}$, finding segments related to the query cue modality:
-$\mathbf{S}_{\text{query}}$ $=\text{TopK}(\text{sim}(q_{embed},\{\mathbf{v}_{k}\mid m_{k}\in\mathcal{M}\}),k)$ (12)
-Second, it defines expanded temporal windows $\mathbf{W}$ around the timestamps $(t_{s,k},t_{e,k})$ of these top-$k$ seed segments, using a temporal buffer $\delta$ to capture adjacent context:
-$\mathbf{W}$ =$=\{[t_{s,k}-\delta,t_{e,k}+\delta]\mid k\in\mathbf{S}_{\text{query}}\}$ (13)
-Third, the system retrieves information from the target modality (complementary to the query cue) by accessing segments $m_{j}$ that temporally overlap with any window in $\mathbf{W}$.
-This leverages temporal proximity for cross-modal association, retrieving, for example, visual details ($\mathbf{T}_{j}^{v},\mathbf{C}_{j}^{v}$) associated with an auditory cue within the same time frame:
-$\mathbf{S}_{\text{target}}$ =$=\{\text{TargetInfo}(m_{j},\text{modality}_{\text{target}})\mid(t_{s,j},t_{e,j})\text{ overlaps with }\mathbf{W}\}$ (14)
-where $\text{TargetInfo}(m_{j},\text{modality}_{\text{target}})$ denotes extracting the specified target modality’s content (e.g., text, embeddings, pointers) from segment $m_{j}$, and the condition $(t_{s,j},t_{e,j})\text{ overlaps with }\mathbf{W}$ is shorthand for overlapping with any interval $[t_{1},t_{2}]\in\mathbf{W}$.
-The hyperparameters $k$ and $\delta$ control the scope and temporal tolerance of this cross-modal retrieval mechanism.
-4.3.3 Adaptive Reasoning for Final Answer Synthesis
-Regardless of the retrieval path (Fast Retrieval or Detailed Recall), the collected evidence $r_{\text{retrieved}}$ must be synthesized into a final answer $a$.
-This is performed by the Adaptive Reasoning module $\Gamma$ (Figure 2b, “Adaptive Reasoning”), corresponding to the “AR” component (Table 1):
-$a=\mathcal{R}_{\text{final}}(q)=\Gamma(q,r_{\text{retrieved}},\mathcal{E}_{\text{context}})$ (15)
-$\Gamma$ is implemented using an LLM that receives query $q$, retrieved evidence $r_{\text{retrieved}}$, and context $\mathcal{E}_{\text{context}}$ with essential visual and auditory information from $r_{\text{retrieved}}$.
-The LLM synthesizes responses by analyzing evidence against the query, prioritizing relevance, resolving inconsistencies, and acknowledging limitations.
-This enables comprehensive responses that go beyond simple aggregation of the retrieved content.
+Adaptive Reasoning.
+While the hippocampus provides memory access, the prefrontal cortex performs executive control and conflict resolution.
+Our Adaptive Reasoning module implements this cognitive function, particularly crucial when reconciling outputs from the dual retrieval pathways.
+The module $\Gamma$ serves as a locus for cognitive reflection, especially when the system escalates from Fast Retrieval to Detailed Recall.
+In such cases, $\Gamma$ receives both summary-based evidence from $\Phi_{\text{fast}}$ and fine-grained evidence from $\Psi_{\text{detailed}}$, reconciling potentially conflicting accounts.
+For instance, if a summary suggests general activity while detailed recall reveals specific contradictory actions, the module prioritizes high-fidelity, grounded evidence from the detailed pathway—mirroring prefrontal-hippocampal interactions where top-down control resolves memory conflicts.
+The module synthesizes a final answer that accurately represents the most reliable evidence:
+$a=\mathcal{R}_{\text{final}}(q)=\Gamma(q,r_{\text{retrieved}},\mathcal{E}_{\text{context}}).$ (12)
+By integrating this reflective step, the architecture ensures its output is not just retrieved facts but a reasoned interpretation, demonstrating that cognitive architectures provide both biological plausibility and functional superiority.
 5 Experiments
 5.1 Implementation Details
-Our system, HippoMM, was implemented on a NVIDIA L40s GPU.
-For Memory Formation, adaptive temporal segmentation processed video frames at 1-10 fps, identifying boundaries using SSIM and audio energy thresholds to create 5-10s segments.
+Our system, HippoMM, was implemented on an NVIDIA L40S GPU.
+Episodic segmentation processed video frames at 1-10 fps using SSIM and audio energy thresholds to create 5-10s segments.
 Perceptual Encoding generated ShortTermMemory objects ($m_{i}$, Eq.
-7) containing 1024-dim ImageBind [girdhar2023imagebind] embeddings ($\mathcal{E}_{c}$), Whisper [radford2023robust] transcriptions ($\mathcal{T}_{a}$), Qwen2.5-VL [Qwen2.5-VL] visual descriptions ($\mathcal{T}_{v}$), and raw data pointers ($\mathbf{C}_{i}$).
-Memory Consolidation filtered segments based on cosine similarity between average ImageBind embeddings ($\mathbf{v}_{i}$) using threshold $\gamma$.
-Semantic Replay employed Qwen2.5-VL as $\phi_{\text{LLM}}$ to generate textual summaries ($\mathbf{S}_{\theta_{k}}$) stored in persistent ThetaEvent objects ($\theta_{k}$).
-For Memory Retrieval, GPT-4o [hurst2024gpt] served as the query classifier ($\mathcal{Q}_{\text{type}}$), directing queries initially to Fast Retrieval ($\Phi_{\text{fast}}$) over ThetaEvent summaries.
-If confidence was below $\tau=0.75$, Detailed Recall ($\Psi_{\text{detailed}}$) was invoked, using ImageBind embeddings for feature search and GPT-4o for semantic search over descriptions ($\mathbf{T}_{k}$).
-GPT-4o was also used as $\Gamma$ for final answer synthesis.
-Additional hyperparameters ($\tau_{v}$, $\tau_{a}$, $k$, $\delta$) are detailed in the Appendix.
+4) containing 1024-dim ImageBind embeddings [9] ($\mathcal{E}_{c}$), Whisper transcriptions [27] ($\mathcal{T}_{a}$), Qwen2.5-VL visual descriptions [2] ($\mathcal{T}_{v}$), and raw data pointers ($\mathbf{C}_{i}$).
+Memory Consolidation filtered segments based on cosine similarity ($\gamma=0.85$) between average ImageBind embeddings.
+Qwen2.5-VL [2] generated textual summaries ($\mathbf{S}_{\theta_{k}}$) for ThetaEvent objects.
+For Hierarchical Memory Retrieval, GPT-4o [15] served as query classifier and final answer synthesizer ($\Gamma$), while Qwen2.5-VL provided confidence scoring during Fast Retrieval ($\tau=0.75$ threshold) and fine-grained visual analysis during Detailed Recall.
+For complete configurations and prompt templates, see Appendix.
 Figure 3: Visualization of consolidated ThetaEvent embedding space and query retrieval via t-SNE projection.
@@
 The semantic summary ($\mathbf{S}_{\theta}$, green star) is central.
-(Right) Text query embeddings (‘dorm’, ‘football’ - stars) retrieve closest caption/transcription embeddings (crosses), linking queries to specific multimodal segments (corresponding frames and text shown).
-5.2 Main Results
-We evaluated HippoMM on the HippoVlog benchmark (introduced in Section 3) designed for long-form audiovisual understanding, comparing it against NotebookLM [^2] https://notebooklm.google.com/ and Video RAG [ren2024videorag].
-These baselines were selected as they represent the only existing systems capable of processing both long-form video and audio content.
-As summarized in Table 1, HippoMM achieves a state-of-the-art average accuracy of 78.2%, significantly outperforming the next-best method (Video RAG) by 14.0%.
-This advantage is particularly evident in visual tasks (66.8% vs.
-Video RAG’s 41.2%), which we attribute to our adaptive memory consolidation preserving salient details more effectively than standard chunking.
-Furthermore, HippoMM demonstrates strong cross-modal reasoning (70.8% A+V accuracy), confirming its ability to capture inter-sensory temporal relationships [teyler2007hippocampal].
-Note that PT and ART metrics are not reported for NotebookLM, as it operates as a web-based service potentially using pre-indexed knowledge, making direct comparison infeasible.
-In terms of efficiency, HippoMM’s adaptive sampling reduces preprocessing time (PT) by 46.2% compared to Video RAG (5.09h vs 9.46h).
-Critically, HippoMM maintains a balanced average response time (ART) of 20.4s.
-In contrast, Video RAG exhibits a much higher ART (112.5s), largely because its retrieval mechanism fails to locate relevant segments in 55.4% of cases, defaulting to costly detailed processing across the entire long video.
-HippoMM’s hierarchical retrieval avoids this pitfall, balancing responsiveness and accuracy.
-5.3 Ablation Studies
-To assess the contribution of individual components, we performed ablation studies by selectively removing Detailed Recall (DR), Fast Retrieval (FR), and Adaptive Reasoning (AR), with results in Table 1 (bottom rows).
-Removing detailed, frame-level analysis (variant ‘w/o DR, AR’) achieved the fastest ART (4.14s) but at the cost of a >6% drop in visual accuracy (to 60.4%), underscoring DR’s role in fine-grained visual understanding.
-Conversely, removing fast, summary-based access (variant ‘w/o FR, AR’) slightly improved cross-modal accuracy (to 72.0%) but drastically increased ART to 27.3s, confirming FR’s effectiveness for rapid gist retrieval.
-Finally, removing only the adaptive reasoning module (‘w/o AR’) resulted in a noticeable drop in overall accuracy (to 76.8%) and semantic understanding (-1.6%), highlighting AR’s function in refining and reconciling information from different retrieval paths.
-These ablations demonstrate that the synergy between rapid summary access (FR), detailed segment analysis (DR), and intelligent result integration (AR) is crucial for HippoMM’s robust performance.
-The experimental results collectively validate our central hypothesis: translating principles inspired by human memory yields a computational framework with significant advantages in accuracy and efficiency for complex, long-form audiovisual understanding tasks.
-5.4 Analysis of HippoMM Mechanisms
-5.4.1 Visualizing Consolidated Event Representations and Retrieval
-HippoMM employs memory consolidation and semantic replay to transform sequences of short-term memories into coherent, semantically rich ThetaEvent representations.
-Figure 3 provides a visualization of one such ThetaEvent, projected into a 2D space using t-SNE based on ImageBind embeddings.
-The left panel shows embeddings from different constituents of the event: visual features (purple squares), auditory features (orange diamonds), derived captions (blue markers), and audio transcriptions (pink circles).
-Notably, the embedding of the generated summary ($\mathbf{S}_{\theta}$, green star) occupies a central position relative to the various modality-specific feature clusters, suggesting its effectiveness in integrating and balancing information across different perceptual inputs.
-The right panel demonstrates the utility of this representation for query-driven retrieval.
-Text queries, such as “dorm” (yellow star query embedding) and “football” (purple star query embedding), are projected into the same space.
-The system identifies the nearest neighbor embeddings within the detailed ShortTermMemory features (marked by yellow and purple crosses respectively).
-For the “dorm” query, the system successfully retrieves a semantically relevant transcription about “stay” which, while not explicitly containing the word “dorm,” correctly identifies the residential setting depicted in the corresponding visual frame.
-For the “football” query, it precisely retrieves a transcription explicitly mentioning the “first football game” and correctly identifies the associated football field visual scene.
-This illustrates how the consolidated ThetaEvent structure supports both abstract summarization and effective retrieval of specific, grounded details from the original multimodal stream.
-5.4.2 Hierarchical Retrieval Dynamics
-Our hierarchical retrieval (Section 4.3) aims to balance speed and accuracy by first attempting fast, summary-based retrieval ($\Phi_{\text{fast}}$) before resorting to detailed segment search ($\Psi_{\text{detailed}}$).
-We analyzed the usage and performance of these pathways across the HippoVlog benchmark queries, visualizing the results in Figure 4.
-This figure displays individual query outcomes plotted against their response times, differentiating pathways by color and correctness by marker style (filled for correct, open for incorrect).
-The analysis revealed that fast retrieval successfully handled 69.1% of queries (visible by the density of points below the 30s threshold), demonstrating the effectiveness of the semantic summaries in ThetaEvent objects for rapid information access.
-Detailed retrieval was primarily invoked for queries requiring fine-grained perceptual information, constituting the remaining 30.9% of cases (points at or above 30s).
-The dashed line marks the 30s threshold.
-As shown in the figure annotations, the two pathways exhibit significant differences in processing efficiency, with fast retrieval achieving an average response time of 6.5s compared to detailed retrieval’s 51.6s.
-This clear bimodal distribution validates our hierarchical design, confirming the system effectively routes queries to the appropriate pathway based on information needs.
-Figure 4: Retrieval pathway analysis with x-axis showing response time.
-Blue markers ($<$30s) and purple markers ($\geq$30s) represent fast and detailed pathways respectively; filled markers indicate correct outcomes, open markers incorrect.
-Dashed line marks the 30s threshold.
-5.5 Generalization Across Audiovisual Tasks
-To assess HippoMM’s generalization capabilities and benchmark its performance against contemporary methods, we established a multifaceted evaluation protocol as detailed in Table 2.
-This protocol integrates metrics from two complementary domains to thoroughly evaluate multimodal understanding.
-First, we assess event representation fidelity through Content Relevance, Semantic Similarity, and Action Alignment metrics, leveraging ground-truth action scripts from the Charades dataset [sigurdsson2016hollywood] (a detailed collection of annotated home activity videos) with evaluations supported by GPT-4o [hurst2024gpt].
-Second, we probe the long-duration temporal understanding of models via the Needle QA accuracy metric, derived from the challenging Needle-in-a-Haystack task in the MLVU benchmark [MLVU], which involves locating precise information within lengthy video sequences ($\geq$ 5 min) through 415 targeted queries.
-The comparison includes leading multimodal systems such as Qwen2.5-Omni [xu2025qwen2] and VideoLlama2 [damonlpsg2024videollama2], alongside GPT-4o serving as a ceiling reference.
-Evaluation constraints include GPT-4o processing only subsampled frames (0.5 fps), and Qwen2.5-Omni’s inability to complete the Needle QA task due to memory constraints.
-Results clearly highlight HippoMM’s strong generalization capabilities.
-Individual metrics are reported along with a consolidated Overall Score, computed as the arithmetic mean after linearly normalizing the original 1–5 scales to a 0–100% range (note that Qwen2.5-Omni’s score excludes Needle QA).
-As presented in Table 2, HippoMM achieves the highest Overall Score (66.9%) among directly comparable systems.
-Crucially, HippoMM demonstrates balanced strengths, excelling in detailed visual metrics such as Action Alignment (top score of 3.14), while significantly outperforming all other evaluated methods on the demanding long-video Needle QA task (73.1%).
-Moreover, HippoMM closely approximates GPT-4o’s ceiling performance, underscoring its robust visual comprehension.
-Collectively, these findings emphasize HippoMM’s exceptional ability to generalize across complex and diverse multimodal contexts.
-Table 2: Comparison of mean evaluation scores.
-R: Relevance, S: Similarity, A: Alignment (Scores 1-5).
-NQA: Needle QA (%), Overall: Overall Score (%).
-GPT-4o is included as a ceiling reference for visual understanding performance.
-Higher values indicate better performance.
-Best results among directly comparable systems are in bold, second best are underlined.
+(Right) Text query embeddings (“dorm,” “football” - stars) retrieve closest caption/transcription embeddings (crosses), linking queries to specific multimodal segments (corresponding frames and text shown).
+5.2 Main Results and Ablation Studies
+HippoMM yields superior accuracy and efficiency.
+Table 1 presents our results on the HippoVlog benchmark, designed to evaluate long-form, cross-modal associative recall.
+HippoMM achieves state-of-the-art 78.2% accuracy, outperforming strong baselines like Ola [21] (74.9%) and VideoLLaMA 3 [41] (73.0%).
+Notably, we implemented a parameter-free cognitive baseline, Simple-Assoc-Memory (SAM), which uses classic Hebbian auto-association on the same episodic features as HippoMM.
+Its poor performance (30.3%) confirms that a naive cognitive mapping is insufficient, validating the necessity of our structured architectural components.
+Furthermore, HippoMM demonstrates superior efficiency: 46.2% reduction in processing time and $5\times$ faster response time than VideoRAG [28], validating that biological structure enables both accuracy and efficiency gains.
+Ablation studies in Table 1 validate each component.
+Removing the Detailed Recall path (HippoMM w/o DR) causes accuracy to drop to 61.2%, confirming its necessity for fine-grained perceptual reasoning.
+Conversely, removing the Fast Retrieval path (HippoMM w/o FR) maintains accuracy (74.6%) but increases average response time to 19.54s, demonstrating the semantic index’s critical role in efficiency.
+A pure Embedding-Only Retrieval (EOR-only) baseline still achieves 71.1%, proving the inherent strength of the underlying memory structure independent of LLM reasoning.
+Finally, replacing GPT-4o with the smaller Qwen2.5-14B maintains strong performance (70.8%), confirming that our cognitive architecture rather than model scale drives effectiveness.
+5.3 Evaluation of Generalization and Core Capabilities
+Episodic segmentation enables better temporal understanding.
+We evaluated HippoMM through a comprehensive protocol spanning event representation and temporal understanding (Table 2).
+For event representation, we measure Content Relevance (R), Semantic Similarity (S), and Action Alignment (A) using Charades [30] with GPT-4o-based evaluation.
+For temporal understanding, we employ the Needle-in-a-Haystack (NQA) task from MLVU [43], which requires locating specific information within videos exceeding five minutes.
+HippoMM achieves 73.1% on NQA, a 46% improvement over VideoLlama 2, while also leading in Semantic Similarity and Action Alignment.
+Qwen2.5-Omni performs well on shorter videos but did not yield NQA results under our evaluation setup.
+These results validate that hippocampus-inspired episodic segmentation enables superior temporal navigation.
+On Video-MME [7] (61.8%) and LongVideoBench [36] (55.2%), HippoMM shows competitive performance; full cross-benchmark comparison is in the Appendix (Table 4).
+Table 2: Evaluation scores across core capabilities.
+R: Relevance, S: Similarity, A: Alignment (1-5 scale).
+NQA: Needle QA (%).
+Overall: Mean normalized score (%).
+GPT-4o included as a strong reference baseline.
+Best results in bold, second best underlined.
 Method R S A NQA (%) Overall (%)
@@
 GPT-4o 3.62 3.44 3.02 64.8 66.6
+5.4 Analysis of HippoMM Mechanisms
+Consolidation enables cross-modal pattern completion.
+Figure 3 visualizes how memory consolidation creates unified ThetaEvent representations.
+The semantic summary ($\mathbf{S}_{\theta}$) occupies a central position among modality-specific clusters, integrating information across perceptual inputs.
+In retrieval, text queries like “dorm” successfully retrieve semantically relevant content (“stay” transcription with residential visuals) despite lexical mismatch, while “football” precisely retrieves “first football game” mentions with field visuals.
+This demonstrates how consolidated structures enable both abstract summarization and specific detail retrieval from partial cross-modal cues.
+Hierarchical Memory Retrieval reduces computational demands.
+Figure 4 analyzes our hierarchical memory retrieval dynamics.
+Fast retrieval successfully handles 69.1% of queries with 6.5s average response time, while detailed retrieval processes complex queries requiring fine-grained perceptual information in 51.6s.
+This 8x speedup for the majority of queries, achieved without sacrificing accuracy, confirms that biological hierarchical structure yields computational efficiency through intelligent query routing.
+Our GPT-4o query classifier achieves 94.1% routing accuracy; even misrouted queries achieve 52.5% accuracy, demonstrating graceful degradation.
+Figure 4: Retrieval pathway analysis with x-axis showing response time.
+Blue markers (<<30s) and purple markers ($\geq$30s) represent fast and detailed pathways respectively; filled markers indicate correct outcomes, open markers incorrect.
+Dashed line marks the 30s threshold.
+Episodic segmentation preserves event integrity.
+We quantify segmentation quality through Event Integrity Rate (EIR), measuring the fraction of ground-truth events fully contained within single segments:
+$\text{EIR}=\frac{1}{|E_{gt}|}\sum_{e\in E_{gt}}\mathbb{I}(\exists s\in S_{gen}:s\text{ fully contains }e)$ (13)
+where $E_{gt}$ denotes ground-truth events, $S_{gen}$ denotes our generated segments, and $\mathbb{I}(\cdot)$ is the indicator function.
+On Charades [30] with ground-truth action annotations, our episodic segmentation achieves 61.6% EIR versus 40.2% for fixed 15-second chunks.
+This improvement demonstrates superior preservation of semantic event structure, crucial for memory formation and query accuracy.
+Consolidation provides noise resilience.
+We evaluated HippoMM’s robustness by introducing varying levels of perceptual noise to the HippoVlog dataset.
+For visual noise, we applied ffmpeg’s noise filter with alls parameters of 10, 25, and 50 for low, medium, and high levels respectively.
+For auditory noise, we mixed white noise at 0.1, 0.3, and 0.5 volume ratios.
+Under visual noise, overall accuracy decreased from the clean baseline of 78.2% to 75.7% (low), 75.0% (medium), and 74.4% (high); under auditory noise, accuracy declined to 77.5%, 77.0%, and 74.7% respectively.
+Even under severe conditions, maximum degradation remains below 4 percentage points.
+This graceful degradation validates that consolidated multimodal representations provide effective redundancy against perceptual corruption, leveraging both pre-trained encoder robustness and the error-correcting properties of cross-modal consolidation.
+Hyperparameter Robustness.
+We validated that our key hyperparameters ($\tau$ and $\tau_{v}$) are empirically optimal.
+The retrieval threshold ($\tau=0.75$) provides a clear speed-accuracy trade-off: lower values needlessly increase ART, while higher values cause accuracy to drop as detailed recall is bypassed.
+Similarly, the visual segmentation threshold ($\tau_{v}=0.65$) is at an optimal point: lower values over-segment and shatter events, while higher values under-segment and incorrectly merge them, with both cases degrading final accuracy.
+Our default values are thus robust and empirically justified.
 6 Conclusion
-We introduced HippoMM, a hippocampus-inspired architecture that translates neurobiological principles into computational mechanisms for multimodal memory.
-Our short-to-long term memory formation process enables efficient consolidation of perceptual details into semantic abstractions.
-While HippoMM outperforms existing methods on audiovisual benchmarks, it struggles with ambiguous event boundaries and complex cross-modal associations under noisy conditions.
-The dual-pathway retrieval system effectively balances efficiency and detail but requires further refinement to handle truly open-domain queries.
-This work demonstrates both the promise and limitations of neuroinspired approaches, revealing that despite progress in biomimetic design, substantial gaps remain between computational implementations and the robust, flexible memory capabilities observed in biological systems.
+HippoMM implements hippocampal principles of pattern separation, associative recall, and memory consolidation for multimodal understanding, achieving 78.2% accuracy on HippoVlog while maintaining 5× faster response than retrieval-augmented baselines.
+Ablations confirm each component’s necessity, while hierarchical retrieval balances efficiency with quality.
+This work demonstrates that cognitive architectures grounded in neuroscience provide principled foundations for multimodal computing systems, suggesting biological memory mechanisms offer promising pathways toward human-like understanding in artificial systems.
+Acknowledgments
+This work was supported in part by NSF CNS-2112562 and ARO W911NF-23-2-0224.
 Appendix
-Appendix A HippoVlog Construction Details
+Appendix A Key Notation
+Table 3: Key notation used throughout the paper
+Symbol Description
+$X=\{x_{t}\}$ Continuous multimodal data stream
+$\mathcal{S}_{t}$ Segmentation trigger indicator at time $t$
+$\text{1}[\cdot]$ Indicator function (returns 1 if true, 0 otherwise)
+$m_{i}$ $i$-th ShortTermMemory object
+$K$ Set of indices for consolidated memories
+$k\in K$ Index in the consolidated set
+$m_{k}$ Consolidated ShortTermMemory object
+$\theta_{k}$ $k$-th ThetaEvent object
+$\mathcal{M}$ Collection of consolidated ShortTermMemory objects
+$\Theta$ Collection of ThetaEvent objects
+$\mathbf{v}_{i},\mathbf{v}_{k}$ Representative embedding for segment $i$ or $k$
+$\mathbf{E}_{i}$ Sequence of cross-modal embeddings for segment $i$
+$\mathbf{T}_{i}$ Aggregated textual outputs for segment $i$
+$\mathbf{C}_{i}$ Pointers to raw audiovisual data for segment $i$
+$\mathbf{S}_{\theta_{k}}$ Semantic summary of ThetaEvent $k$
+$\mathcal{E}_{c}$ Cross-modal embedding model
+$\mathcal{T}_{v}$ Vision-language model for visual descriptions
+$\mathcal{T}_{a}$ Speech recognition model
+$\phi_{\text{LLM}}$ LLM for semantic summary generation
+$\gamma$ Consolidation similarity threshold
+$\tau$ Confidence threshold for retrieval switching
+$\delta$ Temporal window expansion parameter
+$\cos(\cdot,\cdot)$ Cosine similarity function
+$\text{sim}(\cdot,\cdot)$ Similarity function (cosine similarity)
+$\text{TopK}(\cdot,k)$ Function returning $k$ highest-scoring elements
+$\text{FormattedContext}(\cdot)$ Function to structure multimodal content
+$\text{Extract}_{\text{target}}(\cdot)$ Function to extract target modality content
+$(t_{s,i},t_{e,i})$ Start and end times of segment $i$
+$t_{s,i},t_{e,i}]$ Alternative notation for segment boundaries
+$\Phi_{\text{fast}}$ Fast retrieval function
+$\Psi_{\text{detailed}}$ Detailed recall function
+$\Gamma$ Adaptive reasoning module
+$\mathcal{C}_{\text{context}}$ Context information for reasoning
+Appendix B HippoVlog Construction Details
 The HippoVlog dataset was created through a systematic pipeline designed to capture rich multimodal information from daily vlogs and generate challenging memory-based questions.
 The process follows five main stages: video selection, multimodal data sampling, ground truth generation, question generation, and curation.
-A.1 Video Selection and Preprocessing
-  • Source Selection: 25 daily vlogs from YouTube (Creative Commons licensed), with average duration of 15.3 minutes, totaling approximately 682 minutes of content.
+B.1 Video Selection and Preprocessing
+  • Source Selection: 25 daily vlogs from YouTube (Creative Commons licensed), with average duration of 27.3 minutes, totaling approximately 682 minutes of content.
   • Quality Criteria: Videos were chosen based on clear audio and visual quality, natural episodic structure typical of daily life, and diverse activities (cooking, shopping, travel).
   • Format Standardization: All source videos were preprocessed to a standard format (720p resolution at 30fps) with normalized audio levels.
-A.2 Multimodal Data Sampling
+B.2 Multimodal Data Sampling
 To create manageable and representative data points from the full vlogs, we implemented a robust sampling strategy for both audio and video modalities:
@@
     • Metadata about the sampled segments was preserved for future reference.
-A.3 Segment-Level Ground Truth Generation
+B.3 Segment-Level Ground Truth Generation
 Detailed descriptions combining audio and visual information were generated for each sampled 10-second video segment:
@@
     • All content maintained anonymity (using “the person”, “the speaker”)
-A.4 Question Generation
+B.4 Question Generation
 Multiple-choice questions were generated across four categories, with visual questions specifically based on individual frames extracted from video segments:
@@
     • Have exactly one correct answer among four plausible options
-A.5 Question Curation and Finalization
+B.5 Question Curation and Finalization
 The automatically generated questions underwent further curation:
@@
   • Answer Generation: Each question includes one correct answer and three plausible but verifiably incorrect distractors, balanced for length and style.
-A.6 Question Categories and Examples
+B.6 Question Categories and Examples
 Each of the four question categories in HippoVlog targets a specific memory function:
-A.6.1 Cross-Modal Binding (TV×$\times$×A)
+[…]
 Questions requiring integration of information across visual and auditory modalities.
-Example: Cross-Modal Binding Question: “When the main character is sitting in the car wearing a black jacket and white scarf, what does she audibly express?” Options: A: She says she feels excited about the drive.
+Example: Cross-Modal Binding
+Question: “When the main character is sitting in the car wearing a black jacket and white scarf, what does she audibly express?”
+Options:
+  • A: She says she feels excited about the drive.
   • B: She mentions being scared and says it feels weird.
@@
 Explanation: In the video, as the main character sits in the car wearing a black jacket and white scarf, she audibly expresses her nervousness by saying ‘okay, hey, I’m scared, oh my gosh this feels weird.’
-A.6.2 Auditory-Focused Retrieval (TA)
+B.6.2 Auditory-Focused Retrieval (TA)
 Questions focusing on auditory information, including speech content, background sounds, and music.
-Example: Auditory-Focused Retrieval Question: “What did the main character mention about the purpose of the meeting?” Options: A: To finalize a business deal B: To discuss the upcoming project C: To plan a surprise party D: To talk about a vacation Explanation: The main character specifically mentions the ‘upcoming project’ as the purpose of the meeting, indicating a discussion centered around future plans and tasks.
-A.6.3 Visual-Focused Retrieval (TV)
+Example: Auditory-Focused Retrieval
+Question: “What did the main character mention about the purpose of the meeting?”
+Options:
+  • A: To finalize a business deal
+  • B: To discuss the upcoming project
+  • C: To plan a surprise party
+  • D: To talk about a vacation
+Explanation: The main character specifically mentions the ‘upcoming project’ as the purpose of the meeting, indicating a discussion centered around future plans and tasks.
+B.6.3 Visual-Focused Retrieval (TV)
 Questions targeting visual details that require careful observation of scene elements.
-Example: Visual-Focused Retrieval Question: “What distinctive pattern is visible on the roof of the building?” Options: A: Checkerboard B: Chevron C: Striped D: Polka dot Explanation: The roof of the building in the image has a distinctive chevron pattern.
-Example: Visual-Focused Retrieval Question: “What is the main character wearing on her head?” Options: A: A baseball cap B: A knitted hat C: A colorful headscarf D: A headband Explanation: The main character is wearing a colorful headscarf, as seen in the image.
-A.6.4 Semantic/Temporal Reasoning (TS)
+Example: Visual-Focused Retrieval
+Question: “What distinctive pattern is visible on the roof of the building?”
+Options:
+  • A: Checkerboard
+  • B: Chevron
+  • C: Striped
+  • D: Polka dot
+Explanation: The roof of the building in the image has a distinctive chevron pattern.
+Example: Visual-Focused Retrieval
+Question: “What is the main character wearing on her head?”
+Options:
+  • A: A baseball cap
+  • B: A knitted hat
+  • C: A colorful headscarf
+  • D: A headband
+Explanation: The main character is wearing a colorful headscarf, as seen in the image.
+B.6.4 Semantic/Temporal Reasoning (TS)
 Questions requiring integration of information across temporal segments and abstract understanding.
-Example: Semantic/Temporal Reasoning Question: “What was one significant challenge the main character faced during the experiment?” Options: A: Finding time to eat breakfast B: Dealing with brain fog C: Avoiding social media D: Lack of exercise options Explanation: The main character starts the day with a homemade green juice made from various fruits and vegetables.
-Appendix B HippoMM Implementation Details
-This appendix provides the specific implementation choices, models, and parameters used in the HippoMM system, complementing the conceptual description in Section 4.
+Example: Semantic/Temporal Reasoning
+Question: “What was one significant challenge the main character faced during the experiment?”
+Options:
+  • A: Finding time to eat breakfast
+  • B: Dealing with brain fog
+  • C: Avoiding social media
+  • D: Lack of exercise options
+Explanation: Later in the vlog, after the experiment has been underway for several hours, the main character states they are “dealing with a lot of brain fog,” which is affecting their focus on the task.
+Appendix C HippoMM Implementation Details
+This appendix provides the specific implementation choices, models, and parameters used in the HippoMM system, complementing the conceptual description in the Methodology section.
 System Configuration
 All experiments reported were conducted on a single NVIDIA L40S GPU.
-B.1 Memory Formation Implementation
-Temporal Pattern Separation (Sec. 4.2.1)
+C.1 Memory Formation Implementation
+Temporal Pattern Separation
 The process segments continuous video streams based on perceptual changes:
@@
   • Visual Change Detection: A boundary is triggered ($\mathcal{S}_{t}=1$) if the visual dissimilarity between consecutive frames, measured as $d_{v}(F_{t},F_{t-1})=1-\text{SSIM}(F_{t},F_{t-1})$ (Eq.
-5), exceeds a threshold $\tau_{v}=0.65$.
+2), exceeds a threshold $\tau_{v}=0.65$.
   • Auditory Change Detection: A boundary is triggered ($\mathcal{S}_{t}=1$) if the negative log energy of the raw audio signal, $d_{a}(a_{t})$ (Eq.
-6), exceeds a threshold $\tau_{a}=40$.
+3), exceeds a threshold $\tau_{a}=40$.
 This corresponds to the signal amplitude dropping below a threshold (approx.
@@
   • Segment Duration: The resulting segments are constrained to be between $t_{min}=5$ seconds and $t_{max}=10$ seconds long.
-Shorter segments following boundary detection are merged with preceding ones, and longer intervals between boundaries are split to adhere to this constraint, defining the temporal windows $[t_{start,i},t_{end,i}]$.
-Perceptual Encoding (Sec. 4.2.2)
-Each segment $i$ (defined by $[t_{start,i},t_{end,i}]$) is processed to extract multimodal features:
+Shorter segments following boundary detection are merged with preceding ones, and longer intervals between boundaries are split to adhere to this constraint, defining the temporal windows $t_{start,i},t_{end,i}]$.
+Perceptual Encoding
+Each segment $i$ (defined by $t_{start,i},t_{end,i}]$) is processed to extract multimodal features:
   • Cross-Modal Embeddings: We use ImageBind (v1.0) as the joint embedding model ($\mathcal{E}_{c}$) to generate sequences of shared 1024-dimensional embeddings ($\mathbf{E}_{i}$) from visual frames and audio snippets within the segment.
@@
   • ShortTermMemory Object ($m_{i}$): These processed elements are aggregated and stored in a ShortTermMemory object $m_{i}$, as defined in Eq.
-7.
-It encapsulates: pointers to the raw audiovisual snippets ($\mathbf{C}_{i}$), the sequence of ImageBind embeddings ($\mathbf{E}_{i}$), the aggregated Whisper transcriptions and Qwen-VL visual descriptions ($\mathbf{T}_{i}$), segment start/end timestamps ($t_{i},s_{i}$), and related metadata.
-Memory Consolidation (Sec. 4.2.3)
+(4).
+It encapsulates: pointers to the raw audiovisual snippets ($\mathbf{C}_{i}$), the sequence of ImageBind embeddings ($\mathbf{E}_{i}$), the aggregated Whisper transcriptions and Qwen-VL visual descriptions ($\mathbf{T}_{i}$), segment start/end timestamps ($t_{s,i},t_{e,i}$), and related metadata.
+Memory Consolidation
 Redundancy is reduced by filtering similar consecutive segments:
@@
   • Filtering Criterion: A segment $m_{i}$ is retained in the consolidated set $K$ only if the cosine similarity between its representative embedding $\mathbf{v}_{i}$ and the representative embedding $\mathbf{v}_{j}$ of the most recent previously retained segment $m_{j}$ (where $j\in K,j<i$) is below a predefined threshold $\gamma=0.85$, as per Eq.
-8.
+5.
 Segments not meeting this criterion are discarded (or merged).
 The output is the sequence of consolidated objects $\{m_{k}\mid k\in K\}$.
-Semantic Replay (Sec. 4.2.4)
+Semantic Replay
 Abstracted representations are generated for long-term storage:
@@
 It generates a concise textual summary ($\mathbf{S}_{\theta_{k}}$) using a temperature of 0.7 and a maximum length of 300 tokens (see Table 5 for prompt details), as per Eq.
-9.
+6.
   • ThetaEvent Object ($\theta_{k}$): This object constitutes the long-term memory entry, storing the elements defined in Eq.
-10: the representative cross-modal embedding ($\mathbf{v}_{k}$), the generated textual summary ($\mathbf{S}_{\theta_{k}}$), references to the key visual ($\mathbf{I}_{\theta_{k}}$) and auditory ($\mathbf{A}_{\theta_{k}}$) context used, and the segment’s temporal information ($\mathbf{T}_{\theta_{k}}$) which links back to the detailed ShortTermMemory object $m_{k}$.
+7: the representative cross-modal embedding ($\mathbf{v}_{k}$), the generated textual summary ($\mathbf{S}_{\theta_{k}}$), references to the key visual ($\mathbf{I}_{\theta_{k}}$) and auditory ($\mathbf{A}_{\theta_{k}}$) context used, and the segment’s temporal information ($\mathbf{T}_{\theta_{k}}$) which links back to the detailed ShortTermMemory object $m_{k}$.
 The collection $\{\theta_{k}\}$ forms the abstract long-term memory store.
-B.2 Memory Retrieval Implementation
-Query Analysis (Sec. 4.3.1)
+C.2 Memory Retrieval Implementation
+Query Analysis
 User queries are initially processed to determine the retrieval strategy:
@@
 This uses a 3-shot prompt with examples (see Table 6).
-Hierarchical Retrieval (Sec. 4.3.1)
+Hierarchical Retrieval
 Retrieval follows a confidence-gated two-pathway approach (Eq. 8):
-  • Fast Retrieval ($\Phi_{\text{fast}}$): Retrieval is first attempted using the abstract textual summaries ($\mathbf{S}_{\theta_{k}}$) stored in the collection of ThetaEvent objects (=$\mathcal{E}=\{\theta_{k}\}$).
+  • Fast Retrieval ($\Phi_{\text{fast}}$): Retrieval is first attempted using the abstract textual summaries ($\mathbf{S}_{\theta_{k}}$) stored in the collection of ThetaEvent objects ($\mathbf{\Theta}=\{\theta_{k}\}$).
 Confidence scores are determined by internal LLM calibration (Qwen2.5-VL, see Table 7) on a scale of 0-1.
   • Detailed Recall ($\Psi_{\text{detailed}}$): If the confidence score from Fast Retrieval falls below a threshold $\tau=0.75$, retrieval escalates to the Detailed Recall pathway.
-This pathway searches the consolidated ShortTermMemory objects (=$\mathcal{M}=\{m_{k}\mid k\in K\}$), guided by the query type determined by $\mathcal{Q}_{\text{type}}(q)$.
-Detailed Recall Pathways (Sec. 4.3.2)
+This pathway searches the consolidated ShortTermMemory objects ($\mathcal{M}=\{m_{k}\mid k\in K\}$), guided by the query type determined by $\mathcal{Q}_{\text{type}}(q)$.
+Detailed Recall Pathways
 The specific operations within Detailed Recall ($\Psi_{\text{detailed}}$) are guided by the classified query type:
@@
 Text-based search over the aggregated descriptions/transcriptions ($\mathbf{T}_{k}$) serves as a fallback or alternative (Semantic Search).
-See Tables 6 and 7 for related selection prompts.
+See Tables 8 and 9 for related selection prompts.
   • Cross-Modal Association: For Cross-modal query types, retrieval uses temporal co-occurrence.
 First, the top-$k$ seed segments $\mathbf{S}_{\text{query}}$ most similar to $q_{embed}$ (based on $\mathbf{v}_{k}$) are identified (Eq.
-12, with $k=5$).
+9, with $k=5$).
 Second, expanded temporal windows ($\mathbf{W}$) are defined around these seeds using a temporal buffer parameter $\delta=2$ seconds (Eq.
-13).
+10).
 Third, target modality information ($\mathbf{S}_{\text{target}}$) is extracted from any segment $m_{j}$ whose time interval overlaps with any window in $\mathbf{W}$ (Eq.
-14).
+11).
   • Fine-grained Analysis: If necessary for answering the query, specific content within the retrieved ShortTermMemory segments is analyzed further using Qwen2.5-VL (for visual/multimodal questions) or Whisper (for detailed audio content examination).
 This often occurs during the final synthesis stage.
-Adaptive Reasoning (Sec. 4.3.3)
+Adaptive Reasoning
 The final answer is synthesized from the retrieved information:
@@
 It takes the original query ($q$), the retrieved evidence ($r_{\text{retrieved}}$), and relevant extracted context ($\mathcal{E}_{\text{context}}$) as input (Eq.
-15, see Table 9 for prompt).
+12, see Table 11 for prompt).
 For complex reasoning tasks requiring deeper integration or handling potential conflicts (e.g., comparing Fast Retrieval and Detailed Recall outputs), GPT-4o may be used with a 4-shot prompting strategy and temperature of 0.2 (see Table 10 for a related prompt).
-Appendix C Additional Experimental Details
-C.1 Ablation Studies Parameter Settings
-Acronyms refer to components described in Sec. 4: FR (Fast Retrieval, $\Phi_{\text{fast}}$), DR (Detailed Recall, $\Psi_{\text{detailed}}$), AR (Adaptive Reasoning, $\Gamma$).
+Appendix D Cross-Benchmark Generalization
+Table 4: Cross-benchmark generalization comparison.
+Video-MME and LongVideoBench test general comprehension; MLVU-NQA tests precise temporal retrieval, where HippoMM’s episodic memory excels.
+Best results in bold.
+Method Video-MME LongVideoBench MLVU-NQA
+InternVL2 [4] 61.2% 59.3% 48.3%
+VideoLLaMA 3 [41] 66.2% 59.8% 68.3%
+GPT-4o [15] 71.9% 66.7% 64.8%
+HippoMM (Ours) 61.8% 55.2% 73.1%
+Appendix E Additional Experimental Details
+E.1 Ablation Studies Parameter Settings
+Acronyms refer to components described in the Methodology section: FR (Fast Retrieval, $\Phi_{\text{fast}}$), DR (Detailed Recall, $\Psi_{\text{detailed}}$), AR (Adaptive Reasoning, $\Gamma$).
   • HippoMM w/o DR, AR: Uses Fast Retrieval ($\Phi_{\text{fast}}$) only, with direct answer generation from summaries (no Adaptive Reasoning $\Gamma$).
@@
   • HippoMM w/o AR: Uses both retrieval paths ($\Phi_{\text{fast}}$, $\Psi_{\text{detailed}}$) as per the hierarchical logic, but the final output is based on direct extraction or simple aggregation rather than the full synthesis step ($\Gamma$).
-C.2 Performance Metrics
-  • Processing Time (PT): Total time to execute the Memory Formation phase (Sec. 4.2) on raw videos.
-  • Average Response Time (ART): Average time to execute the Memory Retrieval phase (Sec. 4.3) and generate an answer $a$ for a query $q$.
+E.2 Performance Metrics
+  • Processing Time (PT): Total time to execute the Memory Formation phase on raw videos.
+  • Average Response Time (ART): Average time to execute the Memory Retrieval phase and generate an answer $a$ for a query $q$.
   • Accuracy Metrics: Based on strict binary match with human-annotated ground truth answers.
-C.3 Generalization Evaluation
+E.3 Generalization Evaluation
   • Charades Dataset: 100 randomly selected video clips.
@@
   • Overall Score Calculation: Arithmetic mean after linearly normalizing 1-5 scales to 0-100%.
-C.4 Additional Implementation Parameters
+E.4 Additional Implementation Parameters
   • Key Frame Filtering Threshold: Within Perceptual Encoding, a visual similarity threshold (SSIM-based, value 0.3) is used to filter visually redundant frames before description generation to reduce processing load.
@@
 Table 5: Semantic Replay Summarization Prompt ($\phi_{\text{LLM}}$)
-Task Description Please provide a concise one sentence summary of this event based on the video frames descriptions and audio transcription.
-What is happening in this event?
-Input Format Image descriptions: [Concatenated frame captions from key visual context $\mathbf{I}_{\text{visual}}(m_{k})$] Audio transcription: [Concatenated transcribed text from key auditory context $\mathbf{A}_{\text{audio}}(m_{k})$] Output Requirements One sentence summary ($\mathbf{S}_{\theta_{k}}$) capturing the integrated multimodal essence of the event Focus on key actions, subjects, and setting Maintain factual accuracy based on provided descriptions Avoid introducing information not supported by the input
+Task Description Please provide a concise one sentence summary of this event based on the video frames descriptions and audio transcription. What is happening in this event?
+Input Format
+  • Image descriptions: [Concatenated frame captions from key visual context $\mathbf{I}_{\text{visual}}(m_{k})$]
+  • Audio transcription: [Concatenated transcribed text from key auditory context $\mathbf{A}_{\text{audio}}(m_{k})$]
+Output Requirements
+  • One sentence summary ($\mathbf{S}_{\theta_{k}}$) capturing the integrated multimodal essence of the event
+  • Focus on key actions, subjects, and setting
+  • Maintain factual accuracy based on provided descriptions
+  • Avoid introducing information not supported by the input
 Table 6: Query Type Classification Prompt ($\mathcal{Q}_{\text{type}}$)
-Task Description Classify this question into one of these categories based on what type of information is needed to answer it: Classification Categories VIDEO - Questions specifically about visual elements, appearances, or actions that need frame-by-frame analysis, e.g., “what is the main character holding?” (Corresponds to Visual $V$) AUDIO - Questions about sounds, speech, or audio content that need audio analysis, e.g., “what does the main character say/mention?” (Corresponds to Auditory $A$) VIDEO+AUDIO - Questions that require associating both visual and audio information, e.g., “what is the main character doing while saying/mentioning something?” (Corresponds to Cross-modal $V\times A$) SUMMARY - Questions that focus on the overall content or gist of the video, e.g., “what is the main character doing?” (Corresponds to Semantic/Gist) Guidelines If question is about visual details, appearances, or actions, classify as VIDEO If question is about sounds, speech, or audio content, classify as AUDIO If question requires both visual and audio information, classify as VIDEO+AUDIO If question is about the overall content of the video, classify as SUMMARY Input Question: [user query $q$] Output Format Return ONLY one of these exact words: VIDEO, AUDIO, VIDEO+AUDIO, SUMMARY
+Task Description Classify this question into one of these categories based on what type of information is needed to answer it:
+Classification Categories
+  • VIDEO - Questions specifically about visual elements, appearances, or actions that need frame-by-frame analysis, e.g., “what is the main character holding?” (Corresponds to Visual $V$)
+  • AUDIO - Questions about sounds, speech, or audio content that need audio analysis, e.g., “what does the main character say/mention?” (Corresponds to Auditory $A$)
+  • VIDEO+AUDIO - Questions that require associating both visual and audio information, e.g., “what is the main character doing while saying/mentioning something?” (Corresponds to Cross-modal $V\times A$)
+  • SUMMARY - Questions that focus on the overall content or gist of the video, e.g., “what is the main character doing?” (Corresponds to Semantic/Gist)
+Guidelines
+  • If question is about visual details, appearances, or actions, classify as VIDEO
+  • If question is about sounds, speech, or audio content, classify as AUDIO
+  • If question requires both visual and audio information, classify as VIDEO+AUDIO
+  • If question is about the overall content of the video, classify as SUMMARY
+Input Question: [user query $q$]
+Output Format Return ONLY one of these exact words: VIDEO, AUDIO, VIDEO+AUDIO, SUMMARY
 Table 7: Fast Retrieval Attempt Prompt (Part of $\Phi_{\text{fast}}$)
-Task Description Given the following question and relevant event summaries ($\mathbf{S}_{\theta_{k}}$ from $\mathcal{E}$), analyze whether the question can be answered directly or needs specific analysis (triggering Detailed Recall $\Psi_{\text{detailed}}$).
-Input Question: [user query $q$] Event Summaries: [List of relevant event summaries $\mathbf{S}_{\theta_{k}}$ retrieved based on $q$] Guidelines for analysis General questions about overall video content (likely classified as SUMMARY) should be attempted directly from summaries.
+Task Description Given the following question and relevant event summaries ($\mathbf{S}_{\theta_{k}}$ from $\mathbf{\Theta}$), analyze whether the question can be answered directly or needs specific analysis (triggering Detailed Recall $\Psi_{\text{detailed}}$).
+Input Question: [user query $q$]
+Event Summaries: [List of relevant event summaries $\mathbf{S}_{\theta_{k}}$ retrieved based on $q$]
+Guidelines for analysis
+  • General questions about overall video content (likely classified as SUMMARY) should be attempted directly from summaries.
   • Questions about specific visual details (VIDEO), sounds/speech (AUDIO), or combined elements (VIDEO+AUDIO) often require specific analysis beyond summaries.
   • Estimate confidence based on whether summaries contain sufficient information.
-Output Format Must be in one of these two structures: 1.
-If answerable from summaries:
-ANSWER: <your detailed answer based ONLY on summaries>
+Output Format Must be in one of these two structures:
+1. If answerable from summaries:
+ANSWER: <your detailed answer based ONLY on summaries,
+if multiple choice, output one letter>
 CONFIDENCE: <score between 0.0-1.0, e.g., 0.7 or higher if confident>
@@
 (Note: This confidence score is compared against threshold $\tau$)
-Table 6: Visual Segment Selection Prompt (Part of $\Psi_{\text{detailed}}$ Semantic Search)
+Table 8: Visual Segment Selection Prompt (Part of $\Psi_{\text{detailed}}$ - Semantic Search)
 Task Description Given a question (potentially classified as VIDEO or VIDEO+AUDIO), descriptions from consolidated memory segments ($\mathbf{T}_{k}$), and an optional element to search for, identify at most 5 relevant frame indices (within segments $m_{k}$) for answering the question.
-Input Question: [user query $q$] Element to search for: [search query derived from $q$] Frame descriptions (from relevant $\mathbf{T}_{k}$): [Numbered list of frame descriptions/captions] Instructions Return ONLY numbers separated by commas (e.g., “0,3,5,8,12”) corresponding to indices within the provided list.
+Input Question: [user query $q$] Element to search for: [search query derived from $q$]
+Frame descriptions (from relevant $\mathbf{T}_{k}$): [Numbered list of frame descriptions/captions]
+Instructions
+  • Return ONLY numbers separated by commas (e.g., “0,3,5,8,12”) corresponding to indices within the provided list.
   • Return at most 5 indices.
@@
   • If fewer than 5 frames are relevant, return fewer indices.
-Example good responses “0,3,5,8,12” “1,4,7” Output Format Your response (numbers only, comma-separated):
-Table 7: Audio Segment Selection Prompt (Part of $\Psi_{\text{detailed}}$ Semantic Search)
+Example good responses “0,3,5,8,12” “1,4,7”
+Output Format Your response (numbers only, comma-separated):
+Table 9: Audio Segment Selection Prompt (Part of $\Psi_{\text{detailed}}$ - Semantic Search)
 Task Description Given a question (potentially classified as AUDIO or VIDEO+AUDIO) and transcriptions with timestamps from consolidated memory segments ($\mathbf{T}_{k}$), identify the most relevant time frames where the answer might be found.
-Input Question: [user query $q$] Transcriptions (with timestamps from relevant $\mathbf{T}_{k}$): [List of timestamped transcriptions] Instructions Analyze the transcriptions and identify segments most likely to contain the answer.
-Return a JSON array of time frames (relative to the source video).
-Each frame defines a start and end time.
+Input Question: [user query $q$]
+Transcriptions (with timestamps from relevant $\mathbf{T}_{k}$): [List of timestamped transcriptions]
+Instructions
+  • Analyze the transcriptions and identify segments most likely to contain the answer.
+  • Return a JSON array of time frames (relative to the source video). Each frame defines a start and end time.
         {"start": START_TIME, "end": END_TIME},
         {"start": START_TIME, "end": END_TIME}
-Return at most 5 time frames. Include a small buffer around identified time frames (e.g., $\pm$2 seconds, related to $\delta$). If no relevant segments found, return “[]”.
+  • Return at most 5 time frames.
+  • Include a small buffer around identified time frames (e.g., $\pm$2 seconds, related to $\delta$).
+  • If no relevant segments found, return “[]”.
 Output Format Your response (valid JSON only):
@@
 Task Description You potentially have two candidate answers to the same question $q$.
-One derived from Fast Retrieval ($\Phi_{\text{fast}}$, using summaries $\mathcal{E}$) and another from Detailed Recall ($\Psi_{\text{detailed}}$, using detailed segments $\mathcal{M}$).
+One derived from Fast Retrieval ($\Phi_{\text{fast}}$, using summaries $\mathbf{\Theta}$) and another from Detailed Recall ($\Psi_{\text{detailed}}$, using detailed segments $\mathcal{M}$).
 Compare them and provide a final, reconciled answer $a$.
-Input Question: [user query $q$] Answer from summaries (if $\Phi_{\text{fast}}$ produced one): [direct_answer] Confidence from summaries: [direct_confidence] Answer from detailed analysis (if $\Psi_{\text{detailed}}$ was run): [detailed_answer] Supporting Context ($\mathcal{E}_{\text{context}}$): Sample frame captions: [Sample relevant frame captions from $r_{\text{retrieved}}$] Sample transcriptions: [Sample relevant transcriptions from $r_{\text{retrieved}}$] Instructions Compare both answers for consistency, detail, and plausibility given the context.
+Input Question: [user query $q$]
+Answer from summaries (if $\Phi_{\text{fast}}$ produced one): [direct_answer] Confidence from summaries: [direct_confidence]
+Answer from detailed analysis (if $\Psi_{\text{detailed}}$ was run): [detailed_answer]
+Supporting Context ($\mathcal{E}_{\text{context}}$): Sample frame captions: [Sample relevant frame captions from $r_{\text{retrieved}}$] Sample transcriptions: [Sample relevant transcriptions from $r_{\text{retrieved}}$]
+Instructions
+  • Compare both answers for consistency, detail, and plausibility given the context.
   • If they agree, prefer the more detailed/confident answer.
@@
 Task Description Based on the following retrieved context ($r_{\text{retrieved}}$) and extracted evidence ($\mathcal{E}_{\text{context}}$), please synthesize a final answer ($a$) to the question ($q$).
-Input Question: [user query $q$] Overall Video Context (e.g., relevant summaries $\mathbf{S}_{\theta_{k}}$ if available): [List of relevant high-level context] Relevant Retrieved Content (details from $\Phi_{\text{fast}}$ or $\Psi_{\text{detailed}}$): [List of retrieved content descriptions, e.g., frame captions, transcriptions, details from $m_{k}$] Instructions Please provide a clear and specific answer based on the integration of the overall context and the specific retrieved content.
+Input Question: [user query $q$]
+Overall Video Context (e.g., relevant summaries $\mathbf{S}_{\theta_{k}}$ if available): [List of relevant high-level context]
+Relevant Retrieved Content (details from $\Phi_{\text{fast}}$ or $\Psi_{\text{detailed}}$): [List of retrieved content descriptions, e.g., frame captions, transcriptions, details from $m_{k}$]
+Instructions Please provide a clear and specific answer based on the integration of the overall context and the specific retrieved content.
 If the information is insufficient or contradictory, state that or provide the best possible reasoned answer based on the evidence.
 Output should be one letter if the original question was multiple choice.
-Output Format Answer: <Synthesized answer $a$>
+Output Format
+Answer: <Synthesized answer a>